### v1

# Instability of Legendrian knottedness, and non-regular Lagrangian concordances of knots

Abstract.
We show that the family of smoothly non-isotopic Legendrian pretzel knots from [7] that all have the same Legendrian invariants as the standard unknot have front-spuns that are Legendrian isotopic to the front-spun of the unknot. Besides that, we construct the first examples of Lagrangian concordances between Legendrian knots that are not regular, and hence not decomposable. Finally, we show that the relation of Lagrangian concordance between Legendrian knots is not anti-symmetric, and hence does not define a partial order. The latter two results are based upon a new type of flexibility for Lagrangian concordances with stabilised Legendrian ends.
Key words and phrases:
front spinning, non-regular Lagrangian concordance, Legendrian knottedness
2010 Mathematics Subject Classification:
Primary 53D12; Secondary 53D42
1. Introduction and results
1.1. Instability of Legendrian knottedness
The standard Legendrian $\tt{tb}=-1$ unknot $U\subset(\mathbb{R}^{3},dz-y\,dx)$ of the standard contact vector space has zero rotation number and, the standard representative, has a Chekanov–Eliashberg algebra with $\mathbb{Z}[t^{\pm 1}]$-coefficients generated by a single generator $a$ in degree $|a|=1$, and with differential given by $\partial a=1+t$. The first examples of Legendrian knots inside the contact vector-space that are not Legendrian isotopic to $U$, but which have stable-tame isomorphic Chekanov–Eliashberg algebras, were first constructed in [7] by Cornwell–Ng–Sivek. These examples where further explored by Etnyre–Ng in [22, Section 3.4], where it was shown that the Legendrian representatives $\Lambda_{m}$ of the pretzel knots $P(3,-3,-m)$, $m>3$ described in Figure 1 all satisfy this property. Similarly to $U$, the latter Legendrian knots are also Lagrangian slice. However, they are not doubly-slice as shown by Chantraine–Legout [5] and. as a consequence, not Lagrangian cobordant to $U$.
[…]
Figure 1. The front projection of the Legendrian representative $\Lambda_{m}$ of the pretzel knot $P(3,-3,m)$ has $m-3$ crossings in the red box.
Recall that the front spinning construction (sometimes called the Legendrian suspension) is a construction that from a given Legendrian submanifold $\Lambda$ in the standard contact vector space $\mathbb{R}^{2n+1}$ produces the Legendrian embedding $\Sigma_{S^{k}}\Lambda$ of $\Lambda\times S^{k}$ in $\mathbb{R}^{2(k+n)+1}$. This construction for $k=1$ was introduced by Ekholm, Etnyre and Sullivan in [15], and then extended to the case of an arbitrary $k$ by the second author in [26]. This construction admits several extra extensions, see [1, 12, 30, 33]. One can also produce a higher-dimensional Legendrian from a lower-dimensional one by taking the Legendrian product of $\Lambda\subset(Y,\alpha)$ with the exact Lagrangian zero-section $0_{M}\subset T^{*}M$. This gives the Legendrian embedding $\Lambda\times 0_{M}\subset(Y\times T^{*}M,\alpha+p\,dq)$. When $M=S^{n}$ and $Y=\mathbb{R}^{3}$ is the contact vector space, the contact manifold $\mathbb{R}^{3}\times T^{*}S^{n}$ admits a canonical embedding into $\mathbb{R}^{3+2n}$. The front-spun $\Sigma_{S^{n}}\Lambda$ can be identified with the image of the Legendrian product under this embedding. Note that both the above Legendrian product, as well as the front-spun, produces a well-defined Legendrian isotopy class, which only depends of the Legendrian isotopy class of the initial Legendrian $\Lambda$.
Here we investigate the Legendrian product operation and front spinning construction when applied to the above family of knots $\Lambda_{m}$. Since these knots only can be distinguished by smooth topological invariants, while they Legendrian invariants agree, it is natural to expect that their products and spuns all become Legendrian isotopic; see the following two remarks.
Remark 1.1.
When $n\geq 2$, connected $n$-dimensional submanifolds in $(2n+1)$-dimensions that are homotopic are also smoothly isotopic by the classical result of Haefliger [28]. Hence, the Legendrian products and front spuns of $\Lambda_{n}$ and $U$ are smoothly isotopic.
Remark 1.2.
The Legendrian knots $U$ and $\Lambda_{m}$ all have the same classical invariants except smooth knot class, i.e. the same rotation number and Thurston–Bennequin invariant. Hence, their spuns have the same classical invariants by [15, Lemma 4.16] and [26, Lemma 5.1]. Also their Legendrian contact homology DGAs, $\mathcal{A}(\Lambda_{m})$ and $\mathcal{A}(U)$, are stable tame isomorphic for all $m\geq 3$ with arbitrary coefficients. In particular, using the surgery formula from [16], it follows that the partially wrapped Fukaya categories of the two sectors $\mathbb{C}^{2}_{U}$ and $\mathbb{C}^{2}_{\Lambda_{m}}$ are equivalent, where these sectors are obtained by putting stops at the Legendrians $U,\Lambda_{m}\subset\partial_{\infty}\mathbb{C}^{2}$. Since the partially wrapped Fukaya category of a product of sectors satisfies a Künneth-type formula [23], the partially wrapped Fukaya category of the sector $\mathbb{C}^{2}_{\Lambda_{m}}\times T^{*}S^{n}$ with stop $\Lambda_{m}\times 0_{S^{n}}\subset\partial_{\infty}(\mathbb{C}^{2}\times T^{*}S^{n})$ is thus also equivalent to that of $\mathbb{C}^{2}_{U}\times T^{*}S^{n}$ with stop $U\times 0_{S^{n}}\subset\partial_{\infty}(\mathbb{C}^{2}\times T^{*}S^{n})$ whenever $m\geq 3$. Analogous results for the DGA itself should also follow from a version of Künneths formula for Legendrian contact homology from work in progress by Strakoš [36].
Our first result is that these Legendrian product and spuns indeed are Legendrian isotopic.
Theorem 1.3.
$\Sigma_{S^{n}}U$ is Legendrian isotopic to $\Sigma_{S^{n}}\Lambda_{m}$ in $(\mathbb{R}^{2(n+1)+1},dz-\mathbf{y}\,d\mathbf{x})$ for $n\geq 1$ and $m>3$. In fact, something stronger is true: the Legendrian products $U\times 0_{S^{n}}$ and $\Lambda_{m}\times 0_{S^{n}}$ are Legendrian isotopic inside $(\mathbb{R}^{3}\times T^{*}S^{n},dz-y\,dx+p\,dq)$ when $n\geq 1$ and $m>3$.
In Section 1.3 we give two related, but different constructions of the Legendrian isotopy needed for proving Theorem 1.3. First, note that $\Lambda_{m}$, $m>3$, can be obtained by an ambient Legendrian surgery $(\Lambda_{0})_{\eta_{m}}$ performed on $\Lambda_{0}$ consisting of two unlinked Legendrian unknots of $\tt{tb}=-1$ along a Legendrian surgery arc $\eta_{m}$ with boundary $\partial\eta_{m}\subset\Lambda_{0}$; see Figures 18 and 17. The corresponding exact Lagrangian cobordism is described in Figure 2.
[…]
Figure 2. A Lagrangian slice disc filling of $\Lambda_{m}$ that is built from a Lagrangian standard handle-attachment cobordism from two unlinked $\tt{tb}=-1$ unknots to $\Lambda_{m}$, $m>3$. Here $\eta$ is the Legendrian arc along which the corresponding ambient Legendrian surgery is performed.
More precisely, the two crucial properties that we need from this presentation of $\Lambda_{m}$ are as follows:
  1. (1)
    The Legendrian arc $\eta_{m}$ admits a stabilisation in the complement of $\Lambda_{0}$; and
  2. (2)
    The product of the surgery arc $\eta_{m}\times 0_{S^{n}}\subset\mathbb{R}^{3}\times T^{*}S^{n}$ is formally Legendrian isotopic to the spun $\eta_{0}\times 0_{S^{n}}$ of a stabilised Legendrian arc $\eta_{0}$ as shown in Figure 19, for which the Legendrian ambient surgery $(\Lambda_{0})_{\eta_{0}}$ produces the Legendrian standard unknot $U$. Moreover, we need this formal Legendrian isotopy to be supported in the complement of the product $\Lambda_{0}\times 0_{S^{n}}$.
The first strategy uses the fact that the spuns $\eta_{m}\times 0_{S^{n}}$ and $\eta_{0}\times 0_{S^{n}}$ are Legendrian isotopic; this follows from Murphy’s $h$-principle for loose Legendrians [32], which can be applied since spuns of stabilised Legendrians are loose, see [10]. It then follows by construction of Legendrian ambient surgery and the standard Legendrian neighborhood theorem that $\Lambda_{m}\times 0_{S^{n}}$ and $U\times 0_{S^{n}}$ are Legendrian isotopic.
The second strategy utilises the Weinstein handle-body structure on the complement of the Lagrangian handle-attachment cobordism from the unlinked unknots $\Lambda_{0}$ to $\Lambda_{m}$. This is the so-called complementary sectorial cobordism, which is Weinstein since the Lagrangian handle-attachment cobordism is regular in the sense of Eliashberg–Ganatra–Lazarev [18]; see Section 2.1 for the definitions of regularity and complementary cobordism, and Proposition 3.2 for the regularity. Since the handles of the complementary cobordism are attached along stabilised Legendrian knots, the complementary cobordism becomes flexible after spinning. From this we can produce a symplectomorphism of the symplectisation of $\mathbb{R}^{3}\times T^{*}S^{n}$ that identifies the standard Lagrangian fillings of $U\times 0_{S^{n}}$ and $\Lambda_{m}\times 0_{S^{n}}$. The Legendrian isotopy is then constructed by applying Lemma 4.1 to the contactomorphism of the contact boundary that is induced by the symplectomorphism.
1.2. Non-regular Lagrangian concordances in the symplectisation of $\mathbb{R}^{3}$
The existence of a flexible regular fillings is a crucial ingredient for constructing the contact isotopy in Theorem 1.3. While there exist plenty of regular fillings whose Weinstein complements are not flexible, we unfortunately do not have a criterion for the regularity of Lagrangians inside Weinstein cobordisms. It is expected that all exact Lagrangian fillings in Weinstein domains are regular, as formulated by Eliashberg–Ganatra–Lazarev in [18, Problem 2.5]. Furthermore, it was proposed there that any Lagrangian cobordism with a negative end which is not loose or stabilised Legendrian is regular. (In [11] the authors showed that this condition must be weakened to, at least, the condition that the negative end is not sub-loose, which means that it is not cobordant to a loose Legendrian.)
Non-regular Lagrangian cobordisms were first constructed by Eliashberg–Murphy in high dimensions [21] and Lin [31] in dimension four. In these cases, the negative ends are stabilised (which means loose in higher dimensions). These Lagrangian cobordisms are non-regular for the basic topological reason that they have empty positive ends; such a regular oriented Lagrangian $L$ inside a Weinstein cobordism $W$ must define a non-zero homology class $[L]\in H_{n}(\overline{W},\partial_{-}\overline{W})$ relative the concave boundary of $\overline{W}$.
Here we produce the first examples of Lagrangian concordances of Legendrian knots inside the symplectisation of three-dimensional contact vector space that are not regular; see Section 2.1. Note that the negative Legendrian ends of the constructed Lagrangian concordances are stabilised and, hence, they do not admit Lagrangian fillings.
Theorem 1.4 (see Theorem 5.6).
Let $\Lambda$ be a Legendrian knot that admits a decomposable Lagrangian disc filling. Then there exists an exact Lagrangian concordance $C\subset(\mathbb{R}_{t}\times\mathbb{R}^{3},d(e^{t}\alpha_{0}))$ with concave Legendrian boundary $\Lambda_{-}$ obtained by $k$-fold positive and negative stabilisation of $\Lambda$ for $k\gg 0$ sufficiently large, and a convex Legendrian boundary $\Lambda_{+}$ which is the standard Legendrian unknot $U$ stabilised the same number of times.
The proof of Theorem 1.4 relies on recent work by the first author [8], where it is shown that a totally real concordance between Legendrian knots can be approximated by an exact Lagrangian concordance in the same isotopy class, after the knots have been stabilised sufficiently many times with both positive and negative stabilisations. For investigating the totally real concordances, in Appendix A we prove the analogue of Chantraine’s result from [4] on the behaviour of classical invariants for Lagrangian concordances in the case of totally real concordances; see Theorem A.1.
Showing that the constructed Lagrangian concordances are non-regular is more subtle compared to the situation of cobordisms with empty positive ends, and the obstruction that we use is based upon non-trivial input by Kronheimer–Mrowka in gauge theory; in fact, we need the reformulation [7, Theorem 3.1] due to Cornwell–Ng–Sivek. The non-regularity of the concordance follows from our generalisation Theorem 5.6 of [7, Theorem 3.2], where it is shown that a smooth concordance from a non-trivial knot to the unknot must have a local maximum, and can in particular not be decomposable in the sense of Chantraine [4]. In fact, it also follows that the concordance in Theorem 1.4 is not ribbon.
Finally, we show that the non-regular concordances produced by Theorem 1.4 become Hamiltonian isotopic to the standard cylinder over a loose Legendrian torus after spinning; see Remark 5.9.
1.3. Lagrangian concordance of Legendrian knots is not a partial order
We provide negative answer to the reasonably old question of whether the Lagrangian concordance relation between Legendrian knots is a partial order [3, 34]. By definition, this relation $\Lambda_{-}\preceq_{conc}\Lambda_{+}$ holds whenever there exists a Lagrangian concordance with negative and positive end $\Lambda_{-}$ and $\Lambda_{+}$, respectively. The relation is clearly reflexive and transitive, and it has been asked whether the relation defines a partial order. The concordances produced in Theorem 1.4 show that the Lagrangian concordance relation is not anti-symmetric, and hence is not a partial order. More precisely, since any Lagrangian slice disc of a Legendrian $\Lambda_{-}$ induces a Lagrangian concordance from the standard Legendrian unknot $U$ to $\Lambda_{-}$, i.e. $U\preceq_{conc}\Lambda_{-},$ and since the Lagrangian concordance relation can be seen to be preserved under stabilisation of the two knots [3], we get:
Corollary 1.5.
Consider a Legendrian $\Lambda_{-}$ which is obtained by sufficiently many stabilisations (of both signs) of a Legendrian knot that admits filling by a decomposable Lagrangian slice disc, e.g. the non-trivial knot $\overline{9_{46}}$. The Legendrian $\Lambda_{-}$ then admits a concordance both to and from the Legendrian unknot with the same number of stabilisations.
Acknowledgments
The authors are grateful to Baptiste Chantraine, Yakov Eliashberg and Oleg Lazarev for the very helpful discussions. The first author is supported by the Knut and Alice Wallenberg Foundation under the grants KAW 2021.0191 and KAW 2021.0300, and by the Swedish Research Council under the grant number 2020-04426. The second author is supported by the GAČR EXPRO Grant 19-28628X and 4EU+ academic mini-grant MA/4EU+/2024/F3/01.
2. Preliminaries
In this section we give the definition of regular Lagrangian cobordisms, which is due to Eliashberg–Ganatra–Lazarev [18]. Since the complement of a regular Lagrangian cobordism is Weinstein by definition, it can be understood through Kirby calculus, which we also recall.
2.1. Regular Lagrangian cobordisms
A compact Liouville cobordism is a compact exact symplectic cobordism $(\overline{W},d\lambda)$, whose Liouville vector field $\zeta$ is defined uniquely by the equation $d\lambda(\zeta,\cdot)=\lambda(\cdot)$, and which we require to be transverse to the boundary $\partial\overline{W}$; i.e. the boundary is of contact-type, with contact form given by the restriction $\eta|_{T\partial{\overline{W}}}$. By $(\partial_{-}\overline{W},\alpha_{-})$ (resp. $(\partial_{+}\overline{W},\alpha_{+})$) we denote the concave (resp. convex) contact boundary, i.e. the component along which the Liouville vector field points inwards (resp. outwards). Recall that $\partial_{+}\overline{W}\neq\emptyset$ must hold by Stokes’ theorem. The Liouville cobordism is said to be Weinstein if the Liouville vector field is gradient-like for a Morse function.
A compact Lagrangian cobordism in a compact Liouville cobordism is a Lagrangian embedding $(\overline{L},\partial\overline{L})\subset(\overline{W},\partial\overline{W})$ which is cylindrical near the boundary, i.e. tangent to the Liouville vector field there. The decomposition $\partial\overline{W}=\partial_{+}\overline{W}\sqcup\partial_{-}\overline{W}$ into a convex and concave contact boundary induces a decomposition of the boundary $\partial\overline{L}$ of the compact Lagrangian cobordism into a convex and concave Legendrian boundary
±$\partial_{\pm}\overline{L}=\partial\overline{L}\cap\partial_{\pm}\overline{W}\subset(\partial_{\pm}\overline{W},\alpha_{\pm}).$
We will often work with the completed Lagrangian cobordism $L\subset W$ inside the completed Liouville (or Weinstein) cobordism. The completions are constructed by gluing the half-symplectisations
$W=((-\infty,0]_{t}\times\partial_{-}\overline{W},d(e^{t}\alpha_{-}))\>\>\cup\>\>(\overline{W},d\lambda)\>\>\cup\>\>([0,+\infty)_{t}\times\partial_{+}\overline{W},d(e^{t}\alpha_{+}))$
along the boundary of the compact Weinstein cobordism, and gluing cylindrical ends
$L=(-\infty,0]\times\partial_{-}\overline{L}\>\>\cup\>\>\overline{L}\>\>\cup\>\>[0,+\infty)\times\partial_{+}\overline{L}$
to the compact Lagrangian cobordism. A completed Lagrangian (resp. Liouville/Weinstein) cobordism will simply be called a Lagrangian (resp. Liouville/Weinstein) cobordism.
Regular Lagrangian cobordisms inside Weinstein cobordisms were introduced by Eliashberg, Ganatra and Lazarev in [18]. We now recall the definition.
Definition 2.1.
A Lagrangian cobordism $L\subset W$ from $\Lambda_{-}$ to $\Lambda_{+}$ inside a Weinstein cobordism $W$ is regular if the Weinstein structure can be homotoped to one for which the Liouville vector field is tangent to $L$.
The same authors also proved that if a Lagrangian cobordism is regular, then a Weinstein structure exists for which $L$ satisfies the stronger assumptions given in Definition 2.2 below. This stronger type of regularity is also what we mostly will be working with.
Before we give the stronger definition, we recall that the union
$\overline{L}\cup\partial_{-}\overline{W}\subset\overline{W}$
of a compact Lagrangian cobordism and the concave boundary of a Liouville cobordism admits arbitrarily small compact neighborhoods
$\overline{W}_{L}\subset\overline{W}\>\>\text{of}\>\>\overline{L}\cup\partial_{-}\overline{W}\subset\overline{W}$
which, when chosen appropriately, are smooth Weinstein cobordism with concave end $\partial_{-}\overline{W}_{L}=\partial_{-}\overline{W}$, and in which $\overline{L}\subset\overline{W}_{L}$ is regular. Furthermore, the natural Weinstein structure on $D^{*}L$ endows $\overline{W}_{L}$ with a Weinstein structure for which the skeleton of $\overline{W}_{L}$ is entirely contained inside $L$ and hence, in particular, all critical points of the Liouville vector field are contained in $L$. More precisely, in addition to $\zeta$ being tangent to $L$, we can assume that:
  • all critical points of the Liouville vector field $\zeta$ are contained in $L$; and
  • their descending manifolds are entirely contained in $L$.
In other words, the Weinstein handle decomposition of $\overline{W}_{L}$ is induced from a handle-decomposition of $L$.
We are now ready to recall the definition of a special regular Lagrangian from [18].
Definition 2.2.
A Lagrangian cobordism $\overline{L}\subset\overline{W}$ from $\Lambda_{-}\subset\partial_{-}\overline{W}$ to $\Lambda_{+}\subset\partial_{+}\overline{W}$ in a Weinstein cobordism is a special regular Lagrangian if the Weinstein structure on $\overline{W}$ is homotopic to one obtained from $\overline{W}_{L}$ by adding Weinstein handles that are attached away from $\Lambda_{+}\subset\overline{W}_{L}$. We call the induced Weinstein cobordism with concave end $\partial_{+}\overline{W}_{L}$ and convex end $\partial_{+}\overline{W}$ the complementary Weinstein cobordism, and its handles the complementary handles.
Remark 2.3.
Using the language of Weinstein pairs from [17] and Weinstein sectors from [24], the above can be reformulated in the following way. Consider the sector $W_{\Lambda_{+}}$ corresponding to the Weinstein pair $(\overline{W},\Lambda_{+})$, in which $\tilde{L}\coloneqq L\cap W_{\Lambda_{+}}$ is naturally included as a compact Lagrangian with boundary $\partial\tilde{L}$ contained inside the sectorial boundary $\partial W_{\Lambda_{+}}$ of the complete sector. In this setup, $L$ is a special regular Lagrangian iff $\tilde{L}$ can be realised as a subset of the skeleton of $W_{\Lambda_{+}}$ for some homotopic Weinstein structure. In this setting, there is a natural complementary sectorial Weinstein cobordism, which is given as the complement of a neighbourhood of $\tilde{L}$ in the sector.
Since regular Lagrangians have a complement that is Weinstein, they can be presented and studied by using Legendrian handle-body theory, e.g. Kirby diagrams [27] when $\dim W=4$. The regular Lagrangians can also be manipulated via Weinstein homotopy moves which, in this dimension, translate to Kirby moves.
2.2. Standard moves for Kirby diagrams
Here we discuss the tools from Kirby calculus that we need to understand and modify the regular cobordisms that we are interested in. We refer to work by Gompf [27] for a good introduction to Kirby calculus in the setting of contact topology. Our Kirby diagrams will consist of a number of attaching 0-spheres inside $(\mathbb{R}^{3}_{xyz},dz-ydx)$ which have vanishing $y$–coordinates. The Legendrian arcs are allowed to enter the handles both from the left and right in the front projection, where the arcs are required to have constant $z$–coordinate. The attaching 1-spheres will typically be depicted in red colour, while Legendrian knots will be black. (The latter can be thought of as stops of a Weinstein sector that is built by attaching the one and two-handles.)
In order to relate two different diagrams, we need the following moves that relate different Kirby diagrams for the same Weinstein manifold.
Move K–0: Legendrian isotopy in the complement of the attaching spheres: Both Legendrian knots in the contact manifold, as well as attaching 1-spheres, can be deformed by a Legendrian isotopy that supported in the complement of the attaching 0-spheres. By our convention the attaching 0-spheres have a vanishing $y$–coordinate, which means that any Legendrian arc whose front projection has non-zero slope can pass over the small balls where the attaching 0-spheres live. See Figure 3.
Figure 3. The Kirby move K–0: Moving a Legendrian arc of positive slope over a component of the attaching 0-sphere, together with the possible Legendrian arcs entering it.
Move K–I: Introducing/cancelling a pair of handles: A Weinstein one- and two-handle that are in cancelling position can be either introduced or removed. This is a standard Kirby move; see [13, Section 5] in work by Ding and Geiges for the version in contact topology. This move is shown in Figure 4.
Figure 4. The Kirby move K–I: The Kirby move that consists of cancelling a pair consisting of a Weinstein 1-handle and a Weinstein 2-handle in standard position corresponds to the following deformation of the front projection.
Move K–II: Handle-sliding parallel strands: Handling sliding parallel strands of a Legendrian over the two-handle which is in cancelling position gives rise to a move shown in Figure 5. This is a standard Kirby move; the version in contact manifolds was described by Ding and Geiges in [13]; and Casals and Murphy in [2].
Figure 5. The Kirby move K–II: The bottom figure shows the effect in the front projection of the Kirby move that consists of handle-sliding $k=2$ parallel Legendrian strands shown on top over its Legendrian two-handle whose attaching sphere is shown in red.
Move K–III: Introducing a double-stabilisation: Taken any number of parallel strands that pass through a single one-handle, one can move the top one to the bottom, introducing a double stabilisation of the same. See [27, Figure 17] for one version, also its reflection in the z-axis is a possible move. See Figure 6.
Figure 6. The Kirby move K–III: one can cyclically permute the strands going through the handle and introduce a double-stabilization.
Move K–IV: Moving strands around the attaching sphere: The bottom strand to the left (resp. right) of the left (resp. right) attaching sphere can be moved to the top of the right (resp. left) side of the same attaching sphere. The strand is given an additional downward cusp-edge in the direction oriented away from the attaching sphere. See Figure 7.
Figure 7. The Kirby move K–IV: moving a strand entering the one handle from one side of the attaching sphere to the opposite side.
Figure 8. The Kirby move K–IV: moving a strand entering the one handle from one side of the attaching sphere to the opposite side shown in the Lagrangian projection. The top shows the rotation of the arcs performed half-way.
3. Regular Lagrangian cobordisms from Legendrian ambient surgery
The standard technique for constructing a regular Lagrangian cobordisms inside of a symplectisation is by using standard Lagrangian handles that correspond to certain surgery constructions on Legendrians. Notably, in [6] Conway, Etnyre and Tosun showed that decomposable Lagrangian fillings are regular (see Section 3.1 below for the definition). In Proposition 3.2 we generalise this statement to arbitrarily ambient Legendrian surgeries (see Section 3.2 for the definition).
3.1. Decomposable fillings
A particularly nice class of Lagrangian cobordisms in ×$\mathbb{R}\times S^{3}$ are the decomposable ones, which were defined by Chantraine in [4]. By definition, a Lagrangian cobordism is decomposable if it consists of a concatenation of the following simple types of Lagrangian cobordisms:
  • the exact Lagrangian trace-cobordism induced by a Legendrian isotopy;
  • a standard disc-filling of the standard $\tt{tb}=-1$ unknot (this is a zero-handle of the cobordism with an empty negative end), together with other components which all are trivial cylinders; and
  • a Legendrian standard one-handle from a Legendrian knot to the result of a cusp-connected sum performed on it, together with other components which all are trivial cylinders. See Figure 9 for the Legendrians before and after a cusp-connected sum, shown at the bottom and top, respectively.
[…]
Figure 9. Ambient Legendrian surgery: The Legendrian is deformed in a neighborhood of the surgery arc (shown on top) to yield the Legendrian $\Lambda_{\eta}$ on the bottom. There is a standard Lagrangian 1-handle cobordism with concave boundary the Legendrian $\Lambda$ before the surgery, and convex boundary the Legendrian $\Lambda_{\eta}$ produced by the surgery.
[…]
Figure 10. Performing the surgery inside a cancelling handle pair: A small surgery-arc $\eta$ that corresponds to a standard cusp-connected sum of two facing cusp edges can be placed inside a one-handle by a K-II move as shown in this figure. We will indicate the fact that the handle contains the surgery-disc $\eta$ by denoting the cancelling one-handle $S_{\eta}$. The surgery can now be performed inside the handle. cancelling the handle after the surgery by the K-II move again yields the Legendrian $\Lambda_{\eta}$ resulting by the surgery shown in the bottom of Figure 9.
Decomposable Lagrangian cobordisms where shown to be regular in the work of Conway, Etnyre and Tosun [6]. It is still unknown whether the converse is true or not. The proof in [6] moreover gives an explicit recipe for finding the attaching spheres on the complement of the Legendrian. We proceed to explain this recipe in the case of a decomposable Lagrangian filling.
The 0-handles: For each zero-handle we consider $(\mathbb{C}^{2},-d^{c}\rho)$ with the standard filling $\mathfrak{Re}\mathbb{C}^{2}\subset\mathbb{C}^{2}$ of the standard Legendrian unknot $\partial_{\infty}\mathfrak{Re}\mathbb{C}^{2}\subset S^{3}$ of $\tt{tb}=-1$. This filling is clearly is tangent to the radial Liouville vector field, where $\rho(\mathbf{z})=\frac{\|\mathbf{z}\|^{2}}{4}.$ In the general case, when the Lagrangian cobordism has several 0-handles, one has to consider a Weinstein connected sum of such Weinstein four-balls. Since this Weinstein connected sum again produces the standard sphere, and since these handles can be determined from the number of zero-handles, we will omit these 1-handles from the diagram.
The 1-handles: We add a pair of cancelling Weinstein one and two-handles just above the region where the cusp-connected sum takes place; this is move K–I shown in Figure 4. The resulting diagram after having performed the cusp-connected sum is shown at the top of Figure 5. Before the cusp-connected sum, the two facing cusp-edges are now located inside the Weinstein 1-handle. In other words, the cusp-connected sum now takes place entirely inside the one-handle. It can readily be seen that the corresponding standard Lagrangian one-handle in this position is tangent to the standard Liouville structure on the Weinstein one-handle. In particular, the critical point of the Weinstein one-handle coincides with the critical point of the standard Lagrangian one-handle. See Proposition 3.2 for the general statement in arbitrary dimensions.
3.2. The case of a general ambient Legendrian surgery
In [9] the first author generalised the construction of cusp-connected sum to the construction of a Legendrian ambient $k$-surgery on a Legendrian $\Lambda^{n}\subset Y^{2n+1}$, for $k\leq n$. Legendrian ambient $k$-surgery produces an embedding $\Lambda_{\eta}\subset Y$ of the result of a $k$-surgery on $\Lambda$ along a framed sphere $\partial\eta\subset\Lambda$, where $\eta\subset Y$ is a choice of a so-called isotropic surgery $k+1$-disc. The choice of framing of the embedded sphere $\partial\eta$, i.e. a trivialisation of its normal bundle inside $\Lambda$, is also a part of the data, together with a trivialisation of the conformal symplectic normal bundle of $\eta$. The isotropic surgery disc $\eta$ is required to intersect the Legendrian $\Lambda$ precisely along its boundary, where it satisfies the property that the direction in $\eta$ that is normal to $\partial\eta$ also is normal to the Legendrian $\Lambda$. In addition, the trivialisation of the conformal symplectic normal bundle of $\eta$ and the framing of $\partial\eta\subset\Lambda$ are assumed to be compatible in the following sense: the trivialisation of the conformal symplectic normal bundle of $\eta$ is required to be constant along $\partial\eta$ with respect to its framing.
The upshot with the more general definition of Legendrian ambient surgery compared to the cusp-connected sum is that the former construction does not require the Legendrians to be in some particular position; up to contactomorphism the surgery only depends on the choice of Legendrian $\Lambda$, the isotropic surgery disc $\eta$, and the choice of framings of $\partial\eta\subset\Lambda$ and $\eta\subset Y$ up to homotopy.
The Legendrian embedding $\Lambda_{\eta}$ produced by the ambient Legendrian surgery can be assumed to live inside an arbitrarily small neighbourhood of $\Lambda\cup\eta$. In addition, the construction produces an exact Lagrangian handle-attachment cobordism $L_{\eta}$ from $\Lambda$ to $\Lambda_{\eta}$ on which the restriction $t|_{L_{\eta}}$ of the symplectisation coordinate $t$ on $\mathbb{R}_{t}\times Y$ has a unique Morse critical point of index $k+1$.
We refer to [9, Section 4] for the precise description of the construction. The most important feature of the general construction that we rely on here is that the construction is performed by implementing a standard model for the Lagrangian handle in the symplectisation of a standard contact neighbourhood of the isotropic disc $\eta$.
In the case $n=k+1=1$ the construction of the ambient Legendrian surgery in a neighbourhood of $\eta$ is depicted in Figure 11; in this case there is no choice of conformal symplectic normal bundle.
In general, the standard neighbourhood is constructed by using the choice of trivialisation of the conformal symplectic normal bundle of $\eta$, and is of the form
$([-\epsilon,\epsilon]_{z}\times D_{\epsilon}^{*}\eta\times D_{\epsilon}^{*}D^{n-k-1},dz-\lambda_{\eta}-\lambda_{D^{n-k-1}})\hookrightarrow(Y^{2n+1},\alpha),$
where by $\lambda_{M}\in\Omega^{1}(T^{*}M)$ we denote the tautological one-form. More precisely, the surgery $k+1$-disc is identified with the subset
$\{0\}\times 0_{\eta}\times\{0\}\subset\{0\}\times 0_{\eta}\times 0_{D^{n-k-1}}\subset[-\epsilon,\epsilon]_{z}\times D^{*}\eta\times D^{*}D^{n-k-1}$
of the zero-section in the above standard neighbourhood, while the Legendrian $\Lambda$ is identified with the Lagrangian co-normal bundle $N^{*}_{(\partial\eta)\times D^{n-k-1}}=N^{*}_{\partial\eta}\times 0_{D^{n-k-1}}$ of $(\partial\eta)\times D^{n-k-1}$, lifted to have vanishing $[-\epsilon,\epsilon]_{z}$-coordinate.
The choice of conformal symplectic normal bundle induces the identification of the $D^{*}D^{n-k-1}$-factor in the above neighbourhood. Since $N^{*}_{(\partial\eta)\times D^{n-k-1}}$ is a trivial $\mathbb{R}$-fibration over $(\partial\eta)\times D^{n-k-1}$, the compatibility of the framing of $\partial\eta$ and the above conformal symplectic normal bundle implies that the $\mathbb{R}$-fibre of the conormal bundle together with the standard coordinates on the $0_{D^{n-k-1}}$-factor together yield the framing of $\partial\eta\subset\Lambda$.
Passing to the symplectisation, we can find a neighbourhood of $[0,1]\times\eta\subset\mathbb{R}\times Y$ in the symplectisation of the form
$D^{*}[0,1]\times D^{*}\eta\times D^{*}D^{n-k-1}$
in which the cone $[0,1]\times\eta$ is identified with $0_{[0,1]}\times 0_{\eta}\times\{0\}$ and the cone $[0,1]\times\Lambda$ is identified with $0_{[0,1]}\times N^{*}_{\partial\eta}\times 0_{D^{n-k-1}}$. The standard Lagrangian $k+1$-handle is constructed to be a cylinder over $\Lambda$ outside the above neighbourhood of the symplectisation, while the intersection with the above neighbourhood is an explicitly constructed neighbourhood of the core-disc; see [9, Section 4.2.2]. We will only need a description of $L_{\eta}$ in a neighbourhood of the zero-section
$0_{[0,1]}\times 0_{\eta}\times 0_{D^{n-k-1}}\subset D^{*}[0,1]\times D^{*}\eta\times D^{*}D^{n-k-1}.$
Namely, in a small neighbourhood of this zero-section, the standard handle $L_{\eta}$ is given by the Lagrangian co-normal bundle of a $k+1$-dimensional disc in $\Gamma\times\{0\}\subset[0,1]\times\eta\times D^{n-k-1}$ where $\Gamma\subset[0,1]\times\eta$ is a smooth codimension one embedding of a disc given as the union of $[0,1/2]\times\partial\eta$ and a suitable graph of a function $\eta\to[1/2,1]$. In particular, this means that $L_{\eta}$ intersects the zero-section of this standard neighbourhood in a representative $\overline{C}_{\eta}\subset\overline{L}_{\eta}$ of the $k+1$-dimensional core-disc of the handle. Further, we note that the core disc $\overline{C}_{\eta}$ is the boundary of an isotropic $k+2$-disc $\overline{D}_{\eta}\subset 0_{[0,1]}\times 0_{\eta}\times\{0\}$ given by the sub-level set of the above graph. Here $\overline{D}_{\eta}$ is homeomorphic to a disc and has boundary with corners; one boundary stratum is the disc $\overline{C}_{\eta}$, while the other stratum is $\{0\}\times 0_{\eta}\times\{0\}\subset 0_{[0,1]}\times 0_{\eta}\times\{0\}$. By
$C_{\eta}=\overline{C}_{\eta}\cup\left((-\infty,0]\times\partial\eta\right)\>\>\text{and}\>\>D_{\eta}=\overline{D}_{\eta}\cup\left((-\infty,0]\times\partial\eta\right)$
we denote the cylindrical extensions of these discs. In the coordinates used in [9, Section 4.2.2] the $k+1$-disc $C_{\eta}$ is given by the cusp edge of the front projection intersected with $\{x_{k+2}=\ldots=x_{n}=0\}$ while the $k+2$-disc $D_{\eta}$ is the subset of $T^{*}\mathbb{R}^{n+1}$ that is contained in the zero-section intersected with $\{x_{k+2}=\ldots=x_{n}=0\}$, with boundary $C_{\eta}$.
The following lemma is a direct consequence of the construction in terms of a local model implemented in a choice of standard neighbourhood.
Lemma 3.1.
The Legendrian isotopy class of the resulting Legendrian $\Lambda_{\eta}$ only depends on $\Lambda\cup\eta$, together with the data of the framings, up to contact isotopy. The Hamiltonian isotopy class of the resulting Lagrangian standard handle-attachment cobordism
$(\overline{L}_{\eta},\partial\overline{L}_{\eta})\subset([0,1]\times Y,(\{0\}\times\Lambda)\cup(\{1\}\times\Lambda_{\eta}))$
also satisfies the same dependence.
In the special case when $\dim\Lambda=\dim\eta=1$, the Legendrian $\Lambda_{\eta}$ is constructed by performing a cusp-connected sum in a standard neighbourhood $J^{1}\mathbb{R}$ of the arc $\eta$ as outlined in Section 3.3 below.
Using the local model for the surgery we prove the following generalisation of the result from [6] by Conway, Etnyre and Tosun, where they showed that a decomposable Lagrangian cobordism is regular.
Proposition 3.2.
Let $\eta\subset Y^{2n+1}$ be an isotropic surgery disc of dimension $k+1\leq n$ with boundary $\partial\eta\subset\Lambda$. The corresponding standard Lagrangian $k+1$-handle cobordism $L_{\eta}\subset\mathbb{R}\times Y^{2n+1}$ inside the symplectisation is regular; this is a Lagrangian cobordism with concave Legendrian boundary $\Lambda$, and convex boundary the Legendrian $\Lambda_{\eta}$ produced by the surgery. More precisely, the symplectisation admits a Weinstein structure for which:
  1. (1)
    there are two Weinstein handles in cancelling position of index $k+1$ and $k+2$;
  2. (2)
    $L_{\eta}$ is contained in the $k+1$-handle and tangent to the Liouville flow, where the critical point of the Liouville flow of the handle coincides with the critical point of the restriction $t|_{L_{\eta}}$ of the symplectisation coordinate;
  3. (3)
    The descending manifold of the Weinstein $k+1$-handle coincides with the core-disc $C_{\eta}$ consisting of $\overline{C}_{\eta}\subset\overline{L}_{\eta}$ described above with the cylinder $(-\infty,0]\times\partial\eta$ adjoined;
  4. (4)
    The closure of the descending manifold of the Weinstein $k+2$-handle is the isotropic $k+2$-disc $D_{\eta}$ consisting of $\overline{D}_{\eta}$ described above with the cylinder $(-\infty,0]\times\eta$ adjoined.
In the special case when $n=1$ and $k=0$, the corresponding Kirby diagram describing the Legendrian knot $\Lambda_{\eta}$ and the handles can be obtained by replacing the top of Figure 14 (which describes $\Lambda\cup\eta$ in a neighbourhood of $\partial\eta$) with the diagram at the bottom of the same figure.
Proof.
Recall that the standard Weinstein index-$i$ handle of dimension $2(n+1)$ is given by the standard symplectic structure on
$(D^{*}D^{i}\times D^{2(n+1-i)},d(p\,dq)+\omega_{0})$
endowed with a Liouville form for which $\{(0,0)\}$ is a unique critical point of index $i$, whose descending manifold (i.e. the skeleton of the handle, also known as core-disc) is the isotropic $i$-disc $0_{D^{i}}\times\{0\}$, while its ascending manifold (i.e. co-core disc) is the coisotropic $2n-i$-disc $D^{*}_{0}D^{i}\times D^{2(n+1-i)}$. Moreover, the Liouville vector-field can be taken to be the radial Liouville vector field corresponding to the Liouville form $-d^{c}\|\mathbf{z}\|^{2}/4$ in the second component.
We begin by symplectically identifying a neighbourhood of the isotropic $k+1$-disc $C_{\eta}$ with a standard Weinstein $k+1$-handle embedded in ×$\mathbb{R}\times Y$, which has been attached along the isotropic sphere $\partial\eta\subset\partial((-\infty,0]\times Y)$. Here the framing of $\partial\eta$ in $Y$ is induced from the framing of $\partial\eta\subset\Lambda$. Then we denote by $\overline{X}_{a}\subset\mathbb{R}\times Y$ the embedded Weinstein sub-cobordism consisting of $(-\infty,0]\times Y$ together with the attached Weinstein $k+1$-handle. This sub-cobordism has a smooth contact boundary $\partial\overline{X}_{a}\subset\mathbb{R}\times Y$ along which the Liouville vector field induced by the handle-attachment is outwards pointing. (Note that this is different from the original Liouville vector field $\partial_{t}$ on the symplectisation.)
Further, we may assume that this identification takes $L_{\eta}\cap\overline{X}_{a}$ to the Lagrangian
$L_{0}\coloneqq 0_{D^{k+1}}\times(\mathfrak{Im}\mathbb{C}^{n-k}\cap D^{2(n-k)})\subset(D^{*}D^{k+1}\times D^{2(n-k)},d(p\,dq)+\omega_{0})$
contained in the standard Weinstein $k+1$-handle that is everywhere tangent to the Liouville vector field, while $D_{\eta}\cap\overline{X}_{a}$ inside the handle is identified with
$0_{D^{k+1}}\times\left(\{x_{1}\geq 0,x_{2}=\ldots=x_{n-k}=y_{1}=\ldots=y_{n-k}=0\}\cap D^{2(n-k)}\right).$
The intersection of $D_{\eta}$ with the contact boundary $\partial\overline{X}_{a}\subset\mathbb{R}\times Y$ can, moreover, be assumed to be an isotropic $k+1$-sphere.
We may symplectically identify a neighbourhood of $D_{\eta}\setminus X_{a}$ with a Weinstein $k+2$-handle, under which $D_{\eta}\setminus C_{\eta}$ is the isotropic core-disc of the handle. Using $\overline{X}_{b}\supset\overline{X}_{a}$ to denote the union of $\overline{X}_{a}$ together with a neighbourhood of $D_{\eta}$, we have produced a symplectic identification of the neighbourhood $\overline{X}_{b}\subset\mathbb{R}\times Y$ with the result of attaching Weinstein handles of indices $k+1$ and $k+2$ on $\partial(-\infty,0]\times Y$ in cancelling position. In particular, $\partial\overline{X}_{b}$ is contactomorphic to $Y$.
Finally, after a compactly supported deformation of the Liouville vector field $\partial_{t}$ of (×$(\mathbb{R}\times Y,d(e^{t}\alpha))$ to a complete Liouville vector field $\zeta$ without critical points for the same symplectic form, we may assume that $L_{\eta}\setminus X_{b}$ is tangent to $\zeta$ in ×$\mathbb{R}\times Y\setminus\overline{X}_{b}$, and that $\zeta$ coincides with the Liouville vector field induced by the Weinstein handle-decomposition of $\overline{X}_{b}$ near its boundary $\partial\overline{X}_{b}$. This gives us the sought Weinstein structure on ×$\mathbb{R}\times Y$ and thus finishes the proof of properties (1)–(4).
In the special case when $n=1$ and $k=0$, the handles in cancelling position together with the Legendrian $\Lambda_{\eta}$ consists of:
  • A Weinstein one-handle attached at the 0-sphere $\partial\eta$;
  • A Weinstein two-handle attached along a knot $S_{\eta}$ that coincides with $\eta$ outside of the above 00-handle; while
  • The Legendrian $\Lambda_{\eta}$ coincides with $\Lambda$ outside of the one-handle and coincides with a cusp-connected sum of two strands inside the one-handle.
In Darboux balls that contain $\partial\eta$ the Legendrian $\Lambda_{\eta}$ and attaching 00-spheres and $1$-spheres $S_{\eta}$ can be described as shown on the bottom in Figure 13.
Note that this Kirby diagram is not in standard form, since there are strands entering the one-handle from both the left and right. This we correct by the contact isotopy described in K–IV move applied to $\Lambda_{\eta}$; see Figure 7. This contact isotopy is supported inside the one-handle, fixes the attaching sphere $S_{\eta}$ of the cancelling two-handle, and places the strands that correspond to $\Lambda_{\eta}$ on the same side as the strand corresponding to $S_{\eta}$. This produces the configuration shown on the bottom of Figure 14, as sought. ∎
3.3. Ambient surgery, cusp-connected sum, and isotopies.
Since the main focus here is Legendrian ambient 00-surgery on Legendrian knots in $\mathbb{R}^{3}$ we will give some additional details in this case.
When $\eta$ connects two cusp edges in the front projection of the Legendrian link $\Lambda$, the Legendrian ambient surgery can be performed in the following manner.
Step (1): Choose a standard contact jet-neighbourhood $J^{1}\eta$ of the surgery arc in which the Legendrian $\Lambda$ consists of two facing cusp-edges, whose singularities are connected by the zero-section $\eta\subset J^{1}\eta$ that corresponds to the arc $\eta$.
Step (2): Perform a Legendrian isotopy of $\Lambda$ whose support is compact inside the above neighbourhood $J^{1}\eta$, so that the cusp-edges become arbitrarily close to each other; denote the resulting Legendrian by $\Lambda^{\prime}$. We perform this isotopy in an arbitrarily small neighborhood of $\eta$, while fixing the arc $\eta$ set-wise, and so that it induces an isotopy of front-diagrams; See the top and middle front in Figure 11.
Step (3): Perform a standard cusp-connected sum at the two facing cusp-edges of $\Lambda^{\prime}$ that now are joined by a very small segment of $\eta$, and can be assumed to be arbitrarily close. The result is shown in the bottom of Figure 11.
[…]
Figure 11. The top depicts the front projection of the Legendrian link $\Lambda$ and the surgery arc $\eta$ in the standard jet-neighborhood of a surgery-arc. The middle shows the Legendrian $\Lambda^{\prime}$ obtained by a Legendrian isotopy of $\Lambda$ supported near $\eta$ that moves the two cusp-edges closer to each other. The bottom picture depicts the result of a cusp-connected sum of $\Lambda^{\prime}$. This process is the definition of an ambient Legendrian surgery $\Lambda_{\eta}$ of $\Lambda$ defined by the surgery arc $\eta$.
[…]
Figure 12. The Legendrian isotopy of $\Lambda$ following the surgery arc
Moreover, the ambient Legendrian surgery gives rise to a Lagrangian standard handle-attachment cobordism that consists of the trace of the aforementioned isotopy, followed by a Lagrangian standard one-handle.
The isotopy performed in Step (2) is easy to describe also in the front of $\mathbb{R}^{3}$; we express this in a lemma.
Lemma 3.3.
For a generic choice of surgery arc $\eta$ and Legendrian $\Lambda$, the result $\Lambda_{\eta}$ after Legendrian ambient surgery can be obtained by first applying the Legendrian isotopy of $\Lambda$ in Step (2) by performing consecutive moves of type (S1) and (S2) shown in Figure 12 to yield $\Lambda^{\prime}$, and then performing a standard cusp-connected sum on the resulting $\Lambda^{\prime}$.
[…]
Figure 13. Top: The front projection in a Darboux ball that depicts a surgery-arc $\eta$ with boundary on a Legendrian $\Lambda$. Bottom: A Kirby diagram where a one-handle has been attached along $\partial\eta$, a cancelling two-handle has been attached along $S_{\eta}$ which coincides with $\eta$ outside the attaching region, and where $\Lambda_{\eta}$ can be obtained by moving the cusp edges of $\Lambda$ containing $\partial\eta$ towards each other in the one-handle, inside of which a cusp-connected sum is performed. Note that $\Lambda_{\eta}$ and $S_{\eta}$ enter the handles on opposite sides of the attaching sphere.
[…]
Figure 14. The Kirby diagram on the bottom is obtained from the one on the bottom of Figure 13 by a contact isotopy which is the identity in a neighbourhood of $S_{\eta}$, is supported in the one-handle, and which moves $\Lambda_{\eta}$ so that it enters the one-handle on the same side as $S_{\eta}$, with the two strands having larger $z$–coordinates.
4. Spinning surgery discs and attaching spheres (Proof of Theorem 1.3)
In this section we provide two different, but related, strategies that produce Legendrian isotopies after “spinning” a Legendrian $\Lambda\subset Y^{2m+1}$ or, more generally, after passing to the product Legendrian
$\Lambda\times 0_{S^{n}}\subset Y^{2m+1}\times T^{*}S^{n}.$
These strategies will lead to the proof of Theorem 1.3. We will mainly be interested in the case then $Y^{2m+1}=\mathbb{R}^{3}$. We begin by showing some results that can be of general use.
Lemma 4.1.
Assume that $(Y,\xi=\ker\alpha)$ is a contact manifold for which any proper compact subset $K\subsetneq Y$ admits a displacing contact isotopy $\phi_{t}\in\mathrm{Cont}(Y,\xi)$, by which we mean that $\phi_{1}(K)\cap K=\emptyset$. (This is e.g. the case for $Y=\mathbb{R}^{2m+1}$ or the standard contact $S^{2m+1}$.) Then, if $\Lambda\subset(Y\times T^{*}S^{n},\alpha+p\,d\theta)$ is a closed Legendrian, the image $\Psi(\Lambda)$ is Legendrian isotopic to $\Lambda$ whenever $\Psi\in\mathrm{Cont}(Y\times T^{*}S^{n})$ is a contactomorphism with support contained in $K\times T^{*}S^{n}$ for some proper compact subset $K\subsetneq Y$.
Proof.
Any contact isotopy $\phi_{t}\in\mathrm{Cont}(Y,\xi)$ lifts to a contact isotopy
$\Phi_{t}\in\mathrm{Cont}(Y\times T^{*}S^{n}),$
$\Phi_{t}(y,\theta,p)=(\phi_{t}(y),\theta,p\cdot e^{g_{t}(y)})$
where $g_{t}\colon Y\to\mathbb{R}$ is determined by $\phi_{t}^{*}\alpha=e^{g_{t}}\alpha$. For any contactomorphism $\Psi\in\mathrm{Cont}(Y\times T^{*}S^{n})$ supported in $K\times T^{*}S^{n}$ with $K\subsetneq Y$ proper and compact, we can find a contact isotopy of the above form for which $\Phi_{1}$ displaces the support of $\Psi$. The image $\Psi(\Lambda)$ is thus Legendrian isotopic to $\Psi\circ\Phi_{1}(\Lambda)=\Phi_{1}(\Lambda)$ which, in turn, is Legendrian isotopic to $\Phi_{0}(\Lambda)=\Lambda$. ∎
Proposition 4.2.
Assume that two Legendrian links (resp. union of properly embedded arcs) $S_{0},S_{1}\subset Y^{3}\setminus\Lambda$ admit a (compactly supported) Legendrian isotopy $S_{t}\subset Y$, where $\Lambda\subset Y$ is a Legendrian link. It follows that:
  • When $n=2l+1>0$ is odd, the two products $S_{i}\times 0_{S^{n}}$, $i=0,1$, are (compactly supported) formally Legendrian isotopic when considered inside the contact manifold $(Y\times T^{*}S^{n})\setminus(\Lambda\times 0_{S^{n}})$.
  • When $n=2l>0$ is even, the same is true under the additional assumption that any two components of the immersed trace cobordism
    $\{(t,y);\>\>y\in S_{t}\cup\Lambda\}\subset[0,1]\times Y$
    has vanishing algebraic intersection number.
Proof.
In the case when $n$ is odd, we can find an explicitly defined perturbation that removes all intersections of $S_{t}\times 0_{S^{n}}$ and $\Lambda\times 0_{S^{n}}$. More precisely, all these intersections are clean with intersection locus of the form $\{\mathrm{pt}\}\times 0_{S^{n}}$. By the addition of a small non-vanishing section of $T^{*}S^{n}\to S^{n}$ along the intersection loci $\{\mathrm{pt}\}\times 0_{S^{n}}\subset S_{t}\times 0_{S^{n}}$, we obtain a formal Legendrian isotopy from $S_{0}\times 0_{S^{n}}$ to $S_{1}\times 0_{S^{n}}$ that lives inside the complement of $\Lambda\times 0_{S^{n}}$.
In the case when $n$ is even, we argue as follows. The Legendrian isotopy $S_{t}\times 0_{S^{n}}$ intersects $\Lambda\times 0_{S^{n}}$ cleanly in submanifolds of the form $\{\mathrm{pt}\}\times 0_{S^{n}}$. After a generic perturbation of the isotopy, we obtain a situation where the corresponding intersections of the immersed trace cobordism
$\mathcal{T}\coloneqq\{(t,y);\>\>y\in S_{t}\cup\Lambda\}\times 0_{S^{n}}\subset\mathbb{R}\times Y\times T^{*}S^{n}$
are generic double-points where, moreover, the algebraic intersection number of any two fixed components vanishes.
We then deform the isotopy $S_{t}\times 0_{S^{n}}$ through isotopies to one that avoids $\Lambda\times 0_{S}^{n}$ by the following argument using the existence of certain Whitney discs:
Create Whitney disc by connecting two double points of opposite signs with arcs in both intersecting sheets of the immersed trace cobordism $\mathcal{T}$; this forms a loop in $[0,1]_{t}\times Y\times T^{*}S^{n}$. We may assume that each arc intersects each constant $t$-slice $\{t\}\times Y\times T^{*}S^{n}$ in at most one point. Since the Legendrians have codimension strictly greater than two in the contact manifold $Y\times T^{*}S^{n}$, one can find a smooth family of paths in each slice $\{t\}\times Y\times T^{*}S^{n}$ that connect the two unique points on the arcs in the sheets of $\mathcal{T}$. The high codimension makes it possible to assume that these family of paths trace out an embedded two-dimensional disc with boundary on $\mathcal{T}$ and interior disjoint from $\mathcal{T}$. These Whitney discs can be used to deform $S_{t}\times 0_{S^{n}}$ in the required manner, to remove any intersection with $\Lambda\times 0_{S^{n}}$.
Finally, since the initial isotopy was a Legendrian isotopy, the bundle maps can be extended over the family of deformed isotopies that we just constructed, in order to yield the sought formal Legendrian isotopy. ∎
Using the fact that products of stabilised knots become loose [10, Proposition 1.13], combined with Murphy’s $h$-principle for loose Legendrian embeddings [32], we get:
Corollary 4.3.
If the two Legendrians $S_{0},S_{1}\subset Y^{3}\setminus\Lambda$ satisfy the assumptions of Proposition 4.2 and, moreover, are stabilised when considered inside the contact three-manifold $Y^{3}\setminus\Lambda$, then the two Legendrians $S_{i}\times 0_{S^{n}}$, $i=0,1,$ are Legendrian isotopic when considered inside the contact manifold $(Y\times T^{*}S^{n})\setminus(\Lambda\times 0_{S^{n}})$.
Corollary 4.4.
If the two Legendrians $S_{0},S_{1}\subset Y^{3}\setminus\Lambda$ satisfy the assumptions of Proposition 4.2 and, moreover, are stabilised when considered inside the contact three-manifold $Y^{3}\setminus\Lambda$, then the two Legendrians $\Lambda\times 0_{S^{n}}\subset Y_{S_{i}}\times T^{*}S^{n}$, $i=0,1$, are contactomorphic. When $Y_{S_{i}}\cong\mathbb{R}^{3}$, then the induced Legendrians inside $\mathbb{R}^{3}\times T^{*}S^{n}$ are Legendrian isotopic.
We are now ready to prove Theorem 1.3, which follows from either one of the following two constructions.
4.1. First strategy: Spinning the ambient Legendrian surgery discs
We begin with the following basic result.
Lemma 4.5.
Consider a Legendrian $\Lambda\subset Y^{3}$ together with two surgery arcs $\eta_{i}\subset Y$ that have boundary on $\Lambda$ and agree near the boundary. If the two Legendrians
$\eta_{i}\times 0_{S^{n}}\subset Y^{3}\times T^{*}S^{n},\>\>i=0,1,$
are isotopic by a contact isotopy that fixes a neighbourhood of $\Lambda\times 0_{S^{n}}$, then the two spuns $\Lambda_{\eta_{i}}\times 0_{S^{n}}$, $i=0,1,$ are Legendrian isotopic.
Proof.
The ambient Legendrian surgery is constructed in a small standard neighbourhood $J^{1}\eta_{i}$ of $\eta_{i}$ by implementing the Legendrian shown in Figure 11. Since standard neighbourhoods are unique up to contact isotopy, the construction does not depend on the choice of neighbourhood.
Since there is a canonical identification $(J^{1}\eta_{i}\times T^{*}S^{n},dz+p\,dq+\lambda_{S^{n}})=J^{1}(\eta_{i}\times S^{n}),dz+p\,dq+\lambda_{S^{n}})$ of contact manifolds, a product of a contact standard neighbourhood of $\eta_{i}$ with a Weinstein neighbourhood of $0_{S}^{n}$ is a standard neighbourhood of the Legendrian $\eta_{i}\times 0_{S^{n}}$. Any two contact neighbourhoods of $\eta_{i}\times 0_{S^{n}}$, $i=0,1,$ are contact isotopic; here we use the uniqueness of contact neighbourhood together with the assumption that these two Legendrians are isotopic.
It follows that the product Legendrians are isotopic as sought. ∎
Now consider the Legendrian link $\Lambda\subset\mathbb{R}^{3}$ consisting of two unlinked standard unknots, with the surgery arc $\eta_{m}\subset\mathbb{R}^{3}$ as shown in Figure 17 or 18. In particular, the ambient Legendrian surgery produces the Legendrian knot $\Lambda_{\eta_{m}}=\Lambda_{m}$. Then consider the Legendrian surgery arc $\eta_{0}$ for the same Legendrian link $\Lambda$ depicted in Figure 19, for which $\Lambda_{\eta_{0}}$ clearly is Legendrian isotopic to the standard unknot $U$. Note that $\eta_{0}$ and $\eta_{m}$ both are stabilised and Legendrian isotopic in $\mathbb{R}^{3}$ relative boundary. However, any isotopy must clearly intersect $\Lambda$.
In the case when $n$ is odd, we can apply Corollary 4.3 directly to get the sought isotopy between the spun of $\eta_{m}$ and the spun of $\eta_{0}$. Lemma 4.5 then gives the sought Legendrian isotopy between the spuns.
In the case when $n$ is even, we can use the contact isotopy depicted in Figure 15 in order to verify the existence of a Legendrian isotopy that takes the arc $\eta_{m}$ to $\eta_{0}$, and for which the assumptions of Proposition 4.2 are met. Indeed, the obvious isotopy $\eta_{t}$ described in Figure 16 intersects $\Lambda$ twice, but with opposite signs of the corresponding double points on the induced trace cobordism. Again, corollary 4.3 finally shows the claim.
[…]
Figure 15. The configurations showed on the top and second from bottom are contact isotopic.
[…]
Figure 16. The bottom configuration in Figure 15 can be deformed to the top figuration shown here by a Legendrian isotopy of the red arc that intersects $\Lambda_{r}$, i.e. unlinking the red arc from $\Lambda_{r}$. The corresponding immersed trace cobordism has two double points of opposite signs. Finally, after this unlinking, one can apply a global contact isotopy to make both the surgery arcs and Legendrian unknots unlinked as shown at the bottom.
4.2. Second strategy: Spinning the pair of cancelling Weinstein handles
We apply Proposition 3.2 to the standard Lagrangian handle-attachment cobordism whose negative boundary is the two unlinked unknots $\Lambda$ and positive boundary is the Legendrian $\Lambda_{m}$, i.e. the result $\Lambda_{\eta_{m}}$ of an ambient Legendrian surgery on $\Lambda$ along the isotropic arc $\eta_{m}$. The result is a Weinstein structure on ×$\mathbb{R}\times\mathbb{R}^{3}$ consisting of a Weinstein one and two-handle in cancelling position for which the Lagrangian handle-attachment cobordism is tangent to the Liouville flow.
In particular Proposition 3.2 gives us a Legendrian $\tilde{\Lambda}_{\eta}\subset\mathbb{R}^{3}_{sc}$ where the latter contact manifold is the result of a Weinstein 00-surgery on $\mathbb{R}^{3}$, such that $\tilde{\Lambda}_{\eta_{m}}$ becomes $\Lambda_{\eta_{m}}$ after a Weinstein 1-surgery on a knot $S\subset\mathbb{R}^{3}_{sc}\setminus\tilde{\Lambda}_{\eta_{m}}$ that cancels the 00-surgery. The attaching 00-sphere and cancelling $1$-sphere, as well as the knot $\tilde{\Lambda}_{\eta_{m}}$, are obtained from $\Lambda$ and $\eta_{m}$ by the recipe shown in Figure 14.
There exists a stabilised Legendrian knot $S_{0}\subset(\mathbb{R}^{3})_{sc}\setminus\Lambda_{\eta_{m}}$ which also is in cancelling position, but for which $\Lambda_{\eta_{m}}$ becomes the standard unknot $U$ after cancelling the handles. For instance, we can use the ambient Legendrian surgery arc from Figure 19 and replace it with the corresponding attaching sphere from Proposition 3.2; again we can follow the recipe in Figure 14.
Taking products with $T^{*}S^{n}$ results in a generalised Weinstein handle decomposition where $\mathbb{R}^{3}_{sc}\times T^{*}S^{n}$ are obtained from $\mathbb{R}^{3}\times T^{*}S^{n}$ by two subcritical surgeries along isotropic $n$-spheres, and where $\mathbb{R}^{3}\times T^{*}S^{n}$ is obtained from the latter by a generalised Weinstein surgery along the Legendrian $S\times 0_{S^{n}}$; this generalised surgery corresponds to attaching a generalised Weinstein handle of the form $T^{*}(D^{2}\times S^{n})$. Note that such a handle-attachment, as well as the corresponding contact surgery, is completely determined by a choice of Legendrian embedding of $D^{2}\times S^{n}$.
Finally we can apply Corollary 4.4 to the two attaching tori $S_{0}\times 0_{S^{n}}$ and $S\times 0_{S^{n}}$ to deduce the claim that $\Lambda_{m}\times 0_{S^{n}}$ and $U\times 0_{S^{n}}$ are Legendrian isotopic.
Alternatively, one can invoke the $h$-principle for flexible Lagrangian cobordisms [18, Theorem 4.2] to deduce that there is a symplectomorphism of the symplectisation of $\mathbb{R}^{3}\times T^{*}S^{n}$ that identifies the standard Lagrangian fillings of $U\times 0_{S^{n}}$ with the one of $\Lambda_{m}\times 0_{S^{n}}$ (given by the product of the one in Figure 2). The Legendrian isotopy is then constructed by applying Lemma 4.1 to the contactomorphism of the contact boundary that is induced by the symplectomorphism.
[…]
Figure 17. $\Lambda_{m}$, $m>3$ odd, can be produced by an ambient Legendrian surgery along the surgery arc $\eta_{m}$ as shown.
[…]
Figure 18. $\Lambda_{m}$, $m>3$ even, can be produced by an ambient Legendrian surgery along the surgery arc $\eta_{m}$ as shown. See Figure 17 for the first steps.
[…]
Figure 19. Performing the Legendrian ambient surgery on two unlinked unknots, whose front projections moreover are disjoint, along an arc $\eta_{0}$ that intersects the two unknots precisely in their boundary points, produces the standard unknot $U$.
5. Constructing non-regular Lagrangian concordances
The existence of non-regular exact Lagrangian cobordisms were shown by Eliashberg and Murphy in [21] in high dimensions and Lin [31] in dimension four. Namely, these constructions produce exact Lagrangian cobordisms with an empty positive end inside a symplectisation; elementary topological considerations imply that such a cobordism cannot be regular. From this perspective, the example that we construct is different, it has a non-empty positive end and the simplest possible topology among Lagrangian cobordisms, namely that of a concordance.
First we need the following generalisation of [7, Theorem 3.2]
Theorem 5.1.
Let $C\subset[-T,T]\times S^{3}$ be a smooth concordance that coincides with the cylinder over a knot $K_{\pm}$ near the boundary, and such that $[-T,T]\times S^{3}$ can be obtained by attaching $0,1,$ and $2$-handles onto a neighborhood of
($(\{-T\}\times S^{3})\cup C\subset[-T,T]\times S^{3}.$
If $K_{+}$ is the unknot, then so is $K_{-}$.
Proof.
The proof is an adaptation of the proof of [7, Theorem 3.2], which showed that such a concordance $C$ cannot be ribbon, i.e. the symplectisation coordinate must have a local maximum somewhere.
Consider the handle-decomposition of $[-T,T]\times S^{3}$ in which all handles are attached onto a neighbourhood ($(\{-T\}\times S^{3})\cup C$, and for which there are no $3$- or $4$-handles. We may also assume that there are no 00-handles, and that the $2$-handles are attached after the $1$-handles. Since $H_{*}([-T,T]\times S^{3},\{-T\}\times S^{3})=0$, it follows that the number of $1$-handles is equal to the number of $2$-handles; denote this number by $h$.
[…]
Attaching a four-handle to the end $\Sigma(K_{+})=S^{3}$ we obtain a cobordism $X$ from $\Sigma(K_{-})$ to $\emptyset$ which is simply connected. Indeed, turning it upside down, we have a cobordism with only 00-, $2$-, and $3$-handles.
[…]
The rest of the proof is the same as the argument in [7, Theorem 3.2]. In particular, this argument uses the adaptation [7, Theorem 3.1] of a result of Kronheimer and Mrowka that shows that that $\pi_{1}(\Sigma(K_{-}))$ admits a non-trivial representation $\rho$ in $SO(3)$ if and only if $K_{-}$ is not the unknot, while the argument in [29, Section I.3] shows that there can be no such representation under the assumption that $2h$ new generators and relations kill the group $\pi_{1}(\Sigma(K_{-}))$.
[…]
Remark 5.2.
It is unknown to the authors whether a regular Lagrangian cobordism is also decomposable. A decomposable Lagrangian cobordism $L\subset[-T,T]\times S^{3}$ is regular by a particular Weinstein handle decomposition of $[-T,T]\times S^{3}$ in which the handle decomposition of the neighbourhood $T^{*}L$ and the handles attached to $\{-T\}\times S^{3}\cup T^{*}L$ are in cancelling position. It is not clear if a general set of Weinstein handle attachment on $\{-T\}\times S^{3}\cup T^{*}L$ that produces $[-T,T]\times S^{3}$ can be deformed into such a position.
In recent work [8] by the first author it is shown that totally real concordances can be made Lagrangian after adding sufficiently many stabilisations to the Legendrian ends.
Theorem 5.3 ([8]).
Assume that there exists a totally real concordance from a Legendrian knot $\Lambda_{-}^{\prime}$ to $\Lambda_{+}^{\prime}$ (which according to Appendix A implies that $\Lambda_{\pm}^{\prime}$ are smoothly concordant, and with the same ${\tt tb}$ and ${\tt rot}$). Then this concordance can be deformed to a Lagrangian concordance from $\Lambda_{-}$ to $\Lambda_{+}$ by a smooth isotopy, where both $\Lambda_{\pm}$ are obtained from $\Lambda_{\pm}^{\prime}$ by performing a sufficiently large number $k\gg 0$ of stabilisations of both signs.
We will construct totally real concordances in the following general construction, which is based upon the $h$-principle for totally real embeddings.
Proposition 5.4.
If $C\subset\mathbb{R}_{t}\times\mathbb{R}^{3}$ is a decomposable Lagrangian concordance from $\Lambda_{-}$ to $\Lambda_{+}$ with $\tt{rot}(\Lambda_{\pm})=0$, then the reversed cobordism
$C^{\prime}\coloneqq\{(t,x,y,z)\in\mathbb{R}_{t}\times\mathbb{R}^{3};\>(-t,x,y,z)\in C\}$
admits a compactly supported smooth isotopy to a concordance that is totally real for any choice of cylindrical almost complex structure.
Proof.
First, recall that the space of oriented real two-planes $V\subset\mathbb{R}^{4}$ can be identified with $\mathbb{C}P^{1}\times\mathbb{C}P^{1}\cong Gr^{+}_{2}(\mathbb{R}^{4})$. Here we identify the first $\mathbb{C}P^{1}$-factor with an element in the twistor-space $SO(4)/U(2)$, i.e. an almost complex structure $J\in SO(4)$, while the second factor denotes a $J$-complex line in $\mathbb{R}^{4}$. For an appropriate such identification, the Lagrangian planes are given by the subset
$S^{1}\times\mathbb{C}P^{1}\subset\mathbb{C}P^{1}\times\mathbb{C}P^{1}\cong Gr^{+}_{2}(\mathbb{R}^{4}),$
while the non-totally real 2-planes for the standard complex structure $J_{0}$ can be identified with
$D=\{0,\infty\}\times\mathbb{C}P^{1}\subset\mathbb{C}P^{1}\times\mathbb{C}P^{1}\cong Gr^{+}_{2}(\mathbb{R}^{4}).$
More precisely, the latter submanifold consists of those planes that coincide with $J_{0}$-complex planes equipped either with the complex or anti-complex orientation.
It follows from the above that any smooth concordance $C^{\prime}$ between $\Lambda_{\pm}$, where $\tt{rot}(\Lambda_{\pm})=0$, that is cylindrical outside of a compact subset has a Gauss map of tangent planes
$\Gamma\colon C^{\prime}\to Gr^{+}_{2}(\mathbb{R}^{4})$
that is compactly supported homotopic to a family of totally real planes if and only if the intersection number with the hypersurface $D\subset\mathbb{C}P^{1}\times\mathbb{C}P^{1}\cong Gr^{+}_{2}(\mathbb{R}^{4})$ vanishes. Indeed, the assumptions on the boundary imply that the Gauss map along $\partial C^{\prime}$ maps to a null-homologous curve inside the equator $S^{1}\subset\mathbb{C}P^{1}=SO(4)/SU(2)$ when projected to the first factor in the above identification $\mathbb{C}P^{1}\times\mathbb{C}P^{1}\cong Gr^{+}_{2}(\mathbb{R}^{4})$. Thus, the algebraic number of pre-images of the two different components
$\{0\}\times\mathbb{C}P^{1}\>\sqcup\>\{\infty\}\times\mathbb{C}P^{1}=D$
under the Gauss map can readily be seen to agree. When both of these two numbers vanish (which, under the current assumptions, is the same as the vanishing of the total algebraic number of pre-images of $D$), one can readily find a homotopy of the Gauss map relative boundary that removes all intersections with $D$.
By the $h$-principle for totally real embeddings relative to boundary [20, 27.4.1 and Example 2], we simply need to show that $C^{\prime}$ has a Gauss map that is homotopy relative boundary into the totally real planes in order to ensure the existence of the sought smooth isotopy. Using the argument in the previous paragraph, it thus only remains to show that the reversed cobordism $C^{\prime}$ has a Gauss map for which the count of pre-images of $D$ vanishes algebraically.
Since $C$ is decomposable, this follows from a local computation in the standard model. Indeed, by the definition of decomposability, $C$ can be written as a concatenation of Lagrangian traces of Legendrian isotopies and Lagrangian standard 00 and $1$-handles. The subset of $C^{\prime}$ that corresponds to a Lagrangian trace is clearly smoothly isotopic to a totally real cobordism, since it is even smoothly isotopic to the Lagrangian trace of the inverse isotopy. It thus suffices to compute the intersections with $D$ for a standard Lagrangian 00 and $1$-handle. Here, note that away from a small neighbourhood of the critical point of the standard Lagrangian handle, we can assume that the Lagrangian is cylindrical and thus the reversed cobordism is again Lagrangian. Since $C$ is a cylinder there is an equal number of 00 and $1$-handles. It can be seen that these two types of handles contribute with intersection numbers with $D$ of opposite signs; indeed, this can be checked by putting two Lagrangian 00 and $1$-handles in cancelling position. ∎
Remark 5.5.
Note that totally real concordances define a relation which is obviously reflexive and transitive. From Proposition 5.4 it follows that this relation is not anti-symmetric, and hence totally real concordances do not define a partial order.
Theorem 5.6.
Let $\Lambda$ be a Legendrian knot that admits a decomposable Lagrangian disc filling. Then there exists an exact Lagrangian concordance $C\subset(\mathbb{R}_{t}\times\mathbb{R}^{3},d(e^{t}\alpha_{0}))$ with concave Legendrian boundary $\Lambda_{-}$ obtained by $k$-fold positive and negative stabilisation of $\Lambda$ for $k\gg 0$ sufficiently large, and a convex Legendrian boundary $\Lambda_{+}$ which is the standard Legendrian unknot $U$ stabilised the same number of times.
Example 5.7.
Such Lagrangian concordances thus exists from sufficiently stabilised versions of $\Lambda_{m}$, e.g. the $\overline{9_{46}}$-knot $\Lambda_{3}$. See Figures 17 and 18 for the construction of the Lagrangian slice discs.
Remark 5.8.
By Theorem 5.1 the Lagrangian concordances produced by Corollary 5.6 are never regular. Therefore, from the result of Conway, Etnyre and Tosun [6], it follows that these Lagrangian concordances also are not decomposable in the sense of Chantraine [4].
Proof.
A decomposable Lagrangian disc filling can be seen as a standard disc filling of a standard Legendrian unknot $U$ concatenated with a decomposable concordance from $U$ to $\Lambda$. We apply Proposition 5.4 produce a totally real concordance from $\Lambda$ to $U$. Theorem 5.3 then produces the sought Lagrangian concordance, by which we first need to add sufficiently many positive and negative stabilisations of to $U$ and $\Lambda$. ∎
Remark 5.9.
Consider the spun $\Sigma_{S^{1}}C$ of the non-regular concordance $C$ from Corollary 5.6. Note that since by the construction of $C$ the positive and negative ends of $C$ are stabilized Legendrian knots, the positive and negative ends of $\Sigma_{S^{1}}C$ are loose Legendrian tori by [26]. Then using the fact that the positive and negative ends of $\Sigma_{S^{1}}C$ have the same classical Legendrian invariants, they are Legendrian isotopic by the h-principle for loose Legendrian submanifolds, see [32]. From the expected parametric version of h-principle for exact Lagrangian cobordisms with loose negative ends [19] it should follow that $\Sigma_{S^{1}}C$ is Hamiltonian isotopic to the trivial cylinder. Finally, following [26] we see that the same can be said about $\Sigma_{S^{k_{1}}}\dots\Sigma_{S^{k_{r}}}\Sigma_{S^{1}}C$ for $k_{1},\dots,k_{r}\geq 1$.
Appendix A Classical obstructions to the existence of totally real cobordisms between Legendrian knots
Here we prove the direct generalization of the result of Chantraine [3] that provides the classical obstructions to the existence of totally real cobordisms between Legendrian knots.
Theorem A.1.
If there is a totally real orientable cobordism $L$ from Legendrian knot $\Lambda_{-}$ to Legendrian knot $\Lambda_{+}$, then
$tb(\Lambda_{+})-tb(\Lambda_{-})=-\chi(\Lambda),$
$r(\Lambda_{-},[S])=r(\Lambda_{+},[S\cup L]).$
Proof.
The proof is completely analogous to the case when the cobordism is Lagrangian. Namely, we perturb the symplectisation coordinate in a compact subset so that it restricts to a Morse function $h\colon S\to\mathbb{R}$ on the cobordism. Then we consider a push-off by $J(\nabla h)$ for an almost complex structure that is cylindrical outside of a compact subset, compatible with the symplectic structure, and for which the cobordism is totally-real. We can assume that $\nabla h=\partial_{t}$ is satisfied outside of a compact subset. It follows that the signed count of intersection points are precisely $-\chi(S)$ (as in the Lagrangian case). The rest follows in the same manner. (Recall that the Thurston-Bennequin invariant is the self-intersection number computed by the push-off of a Legendrian along the vector field $J\partial_{t}$.) ∎

### v2

@@
 1.1. Instability of Legendrian knottedness
-The standard Legendrian $\tt{tb}=-1$ unknot $U\subset(\mathbb{R}^{3},dz-y\,dx)$ of the standard contact vector space has zero rotation number and, the standard representative, has a Chekanov–Eliashberg algebra with $\mathbb{Z}[t^{\pm 1}]$-coefficients generated by a single generator $a$ in degree $|a|=1$, and with differential given by $\partial a=1+t$.
+In this article we investigate a set of Legendrian knots in $(\mathbb{R}^{3},dz-y\,dx)$ that are not smoothly isotopic, admit exact Lagrangian fillings, and have stable-tame isomorphic Chekanov–Eliashberg algebras.
+(The latter DGAs are not acyclic in view of the fillability.) Recall that the standard Legendrian $\tt{tb}=-1$ unknot $U$ in the standard contact $3$-dimensional vector space has zero rotation number and a Chekanov–Eliashberg algebra with $\mathbb{Z}[t^{\pm 1}]$-coefficients generated by a single generator $a$ in degree $|a|=1$, and with differential given by $\partial a=1+t$.
 The first examples of Legendrian knots inside the contact vector-space that are not Legendrian isotopic to $U$, but which have stable-tame isomorphic Chekanov–Eliashberg algebras, were first constructed in [7] by Cornwell–Ng–Sivek.
-These examples where further explored by Etnyre–Ng in [22, Section 3.4], where it was shown that the Legendrian representatives $\Lambda_{m}$ of the pretzel knots $P(3,-3,-m)$, $m>3$ described in Figure 1 all satisfy this property.
+These examples where further explored by Etnyre–Ng in [23, Section 3.4], where it was shown that the Legendrian representatives $\Lambda_{m}$ of the pretzel knots P(3,3,m)P(3,-3,-m), $m>3$ described in Figure 1 all satisfy this property.
 Similarly to $U$, the latter Legendrian knots are also Lagrangian slice.
-However, they are not doubly-slice as shown by Chantraine–Legout [5] and.
-as a consequence, not Lagrangian cobordant to $U$.
+However, they are not doubly-slice as shown by Chantraine–Legout [5] and, as a consequence, not Lagrangian cobordant to $U$.
 […]
@@
 […]
-Figure 1. The front projection of the Legendrian representative $\Lambda_{m}$ of the pretzel knot $P(3,-3,m)$ has $m-3$ crossings in the red box.
+Figure 1. The front projection of the Legendrian representative $\Lambda_{m}$ of the pretzel knot P(3,3,m)P(3,-3,m) has m3m-3 crossings in the red box.
 Recall that the front spinning construction (sometimes called the Legendrian suspension) is a construction that from a given Legendrian submanifold $\Lambda$ in the standard contact vector space $\mathbb{R}^{2n+1}$ produces the Legendrian embedding $\Sigma_{S^{k}}\Lambda$ of $\Lambda\times S^{k}$ in $\mathbb{R}^{2(k+n)+1}$.
 This construction for $k=1$ was introduced by Ekholm, Etnyre and Sullivan in [16], and then extended to the case of an arbitrary $k$ by the second author in [27].
-This construction admits several extra extensions, see [1, 12, 30, 33].
+This construction admits several generalisations in different directions, see [1, 13, 32, 36].
 One can also produce a higher-dimensional Legendrian from a lower-dimensional one by taking the Legendrian product of $\Lambda\subset(Y,\alpha)$ with the exact Lagrangian zero-section $0_{M}\subset T^{*}M$.
@@
 Here we investigate the Legendrian product operation and front spinning construction when applied to the above family of knots $\Lambda_{m}$.
-Since these knots only can be distinguished by smooth topological invariants, while they Legendrian invariants agree, it is natural to expect that their products and spuns all become Legendrian isotopic; see the following two remarks.
+Since these knots only can be distinguished by smooth topological invariants, while the Legendrian invariants agree, it is natural to suspect that their products and spuns all become Legendrian isotopic; see the following two remarks.
 Remark 1.1.
@@
 the same rotation number and Thurston–Bennequin invariant.
-Hence, their spuns have the same classical invariants by [15, Lemma 4.16] and [26, Lemma 5.1].
+Hence, their spuns have the same classical invariants by [16, Lemma 4.16] and [27, Lemma 5.1].
 Also their Legendrian contact homology DGAs, $\mathcal{A}(\Lambda_{m})$ and $\mathcal{A}(U)$, are stable tame isomorphic for all $m\geq 3$ with arbitrary coefficients.
@@
 In fact, something stronger is true: the Legendrian products $U\times 0_{S^{n}}$ and $\Lambda_{m}\times 0_{S^{n}}$ are Legendrian isotopic inside $(\mathbb{R}^{3}\times T^{*}S^{n},dz-y\,dx+p\,dq)$ when $n\geq 1$ and $m>3$.
-In Section 1.3 we give two related, but different constructions of the Legendrian isotopy needed for proving Theorem 1.3.
+In Section 4 we give two related, but different constructions of the Legendrian isotopy needed for proving Theorem 1.3.
 First, note that $\Lambda_{m}$, $m>3$, can be obtained by an ambient Legendrian surgery $(\Lambda_{0})_{\eta_{m}}$ performed on $\Lambda_{0}$ consisting of two unlinked Legendrian unknots of $\tt{tb}=-1$ along a Legendrian surgery arc $\eta_{m}$ with boundary $\partial\eta_{m}\subset\Lambda_{0}$; see Figures 18 and 17.
@@
   2. (2)
-    The product of the surgery arc $\eta_{m}\times 0_{S^{n}}\subset\mathbb{R}^{3}\times T^{*}S^{n}$ is formally Legendrian isotopic to the spun $\eta_{0}\times 0_{S^{n}}$ of a stabilised Legendrian arc $\eta_{0}$ as shown in Figure 19, for which the Legendrian ambient surgery $(\Lambda_{0})_{\eta_{0}}$ produces the Legendrian standard unknot $U$.
+    The product of the surgery arc $\eta_{m}\times 0_{S^{n}}\subset\mathbb{R}^{3}\times T^{*}S^{n}$ is formally Legendrian isotopic to the spun $\eta^{0}_{m}\times 0_{S^{n}}$ of a stabilised Legendrian arc $\eta^{0}_{m}$ as shown in Figure 19, for which the Legendrian ambient surgery $(\Lambda_{0})_{\eta^{0}_{m}}$ produces the Legendrian standard unknot $U$.
 Moreover, we need this formal Legendrian isotopy to be supported in the complement of the product $\Lambda_{0}\times 0_{S^{n}}$.
-The first strategy uses the fact that the spuns $\eta_{m}\times 0_{S^{n}}$ and $\eta_{0}\times 0_{S^{n}}$ are Legendrian isotopic; this follows from Murphy’s $h$-principle for loose Legendrians [32], which can be applied since spuns of stabilised Legendrians are loose, see [10].
+The first strategy uses the fact that the spuns $\eta_{m}\times 0_{S^{n}}$ and $\eta^{0}_{m}\times 0_{S^{n}}$ are Legendrian isotopic by an isotopy supported in the complement of $\Lambda_{0}\times 0_{S^{n}}$; this follows from Murphy’s $h$-principle for loose Legendrians [34], which can be applied since spuns of stabilised Legendrians are loose, see [11].
 It then follows by construction of Legendrian ambient surgery and the standard Legendrian neighborhood theorem that $\Lambda_{m}\times 0_{S^{n}}$ and $U\times 0_{S^{n}}$ are Legendrian isotopic.
@@
 1.2. Non-regular Lagrangian concordances in the symplectisation of $\mathbb{R}^{3}$
-The existence of a flexible regular fillings is a crucial ingredient for constructing the contact isotopy in Theorem 1.3.
-While there exist plenty of regular fillings whose Weinstein complements are not flexible, we unfortunately do not have a criterion for the regularity of Lagrangians inside Weinstein cobordisms.
-It is expected that all exact Lagrangian fillings in Weinstein domains are regular, as formulated by Eliashberg–Ganatra–Lazarev in [18, Problem 2.5].
+The existence of a flexible regular filling is a crucial ingredient for constructing the contact isotopy in Theorem 1.3.
+While there exist plenty of regular fillings whose Weinstein complements are not flexible, we unfortunately do not have a good understanding of when regularity holds for Lagrangians inside Weinstein cobordisms.
+It is expected that all exact Lagrangian fillings in Weinstein domains are regular, as formulated by Eliashberg–Ganatra–Lazarev in [19, Problem 2.5].
 Furthermore, it was proposed there that any Lagrangian cobordism with a negative end which is not loose or stabilised Legendrian is regular.
@@
 In these cases, the negative ends are stabilised (which means loose in higher dimensions).
-These Lagrangian cobordisms are non-regular for the basic topological reason that they have empty positive ends; such a regular oriented Lagrangian $L$ inside a Weinstein cobordism $W$ must define a non-zero homology class $[L]\in H_{n}(\overline{W},\partial_{-}\overline{W})$ relative the concave boundary of $\overline{W}$.
+These Lagrangian cobordisms are non-regular for the basic topological reason that they have empty positive ends; a regular oriented Lagrangian $L$ inside a Weinstein cobordism $W$ must define a non-zero homology class $L]\in H_{n}(\overline{W},\partial_{-}\overline{W})$ relative the concave boundary of $\overline{W}$.
 Here we produce the first examples of Lagrangian concordances of Legendrian knots inside the symplectisation of three-dimensional contact vector space that are not regular; see Section 2.1.
@@
 1.3. Lagrangian concordance of Legendrian knots is not a partial order
-We provide negative answer to the reasonably old question of whether the Lagrangian concordance relation between Legendrian knots is a partial order [3, 34].
+We provide negative answer to the reasonably old question of whether the Lagrangian concordance relation between Legendrian isotopy classes of Legendrian knots is a partial order [3, 37].
 By definition, this relation $\Lambda_{-}\preceq_{conc}\Lambda_{+}$ holds whenever there exists a Lagrangian concordance with negative and positive end $\Lambda_{-}$ and $\Lambda_{+}$, respectively.
@@
 A compact Liouville cobordism is a compact exact symplectic cobordism $(\overline{W},d\lambda)$, whose Liouville vector field $\zeta$ is defined uniquely by the equation $d\lambda(\zeta,\cdot)=\lambda(\cdot)$, and which we require to be transverse to the boundary $\partial\overline{W}$; i.e.
-the boundary is of contact-type, with contact form given by the restriction $\eta|_{T\partial{\overline{W}}}$.
+the boundary is of contact-type, with contact form given by the restriction $\lambda|_{T\partial{\overline{W}}}$.
 By $(\partial_{-}\overline{W},\alpha_{-})$ (resp.
@@
 The decomposition $\partial\overline{W}=\partial_{+}\overline{W}\sqcup\partial_{-}\overline{W}$ into a convex and concave contact boundary induces a decomposition of the boundary $\partial\overline{L}$ of the compact Lagrangian cobordism into a convex and concave Legendrian boundary
-±$\partial_{\pm}\overline{L}=\partial\overline{L}\cap\partial_{\pm}\overline{W}\subset(\partial_{\pm}\overline{W},\alpha_{\pm}).$
+$\partial_{\pm}\overline{L}=\partial\overline{L}\cap\partial_{\pm}\overline{W}\subset(\partial_{\pm}\overline{W},\alpha_{\pm}).$
 We will often work with the completed Lagrangian cobordism $L\subset W$ inside the completed Liouville (or Weinstein) cobordism. The completions are constructed by gluing the half-symplectisations
@@
 This stronger type of regularity is also what we mostly will be working with.
-Before we give the stronger definition, we recall that the union
+Before we give the stronger definition, we recall, following the discussion in [19, Section 2], that the union
 $\overline{L}\cup\partial_{-}\overline{W}\subset\overline{W}$
@@
 $\overline{W}_{L}\subset\overline{W}\>\>\text{of}\>\>\overline{L}\cup\partial_{-}\overline{W}\subset\overline{W}$
-which, when chosen appropriately, are smooth Weinstein cobordism with concave end $\partial_{-}\overline{W}_{L}=\partial_{-}\overline{W}$, and in which $\overline{L}\subset\overline{W}_{L}$ is regular.
+which, when chosen appropriately, is a smooth compact Weinstein cobordism with concave end $\partial_{-}\overline{W}_{L}=\partial_{-}\overline{W}$, and in which $\overline{L}\subset\overline{W}_{L}$ is regular.
 Furthermore, the natural Weinstein structure on $D^{*}L$ endows $\overline{W}_{L}$ with a Weinstein structure for which the skeleton of $\overline{W}_{L}$ is entirely contained inside $L$ and hence, in particular, all critical points of the Liouville vector field are contained in $L$.
@@
 Consider the sector $W_{\Lambda_{+}}$ corresponding to the Weinstein pair $(\overline{W},\Lambda_{+})$, in which $\tilde{L}\coloneqq L\cap W_{\Lambda_{+}}$ is naturally included as a compact Lagrangian with boundary $\partial\tilde{L}$ contained inside the sectorial boundary $\partial W_{\Lambda_{+}}$ of the complete sector.
-In this setup, $L$ is a special regular Lagrangian iff $\tilde{L}$ can be realised as a subset of the skeleton of $W_{\Lambda_{+}}$ for some homotopic Weinstein structure.
+In this setup, $L$ is a special regular Lagrangian if and only if $\tilde{L}$ can be realised as a subset of the skeleton of $W_{\Lambda_{+}}$ for some homotopic Weinstein structure.
 In this setting, there is a natural complementary sectorial Weinstein cobordism, which is given as the complement of a neighbourhood of $\tilde{L}$ in the sector.
@@
 We refer to work by Gompf [28] for a good introduction to Kirby calculus in the setting of contact topology.
-Our Kirby diagrams will consist of a number of attaching 0-spheres inside $(\mathbb{R}^{3}_{xyz},dz-ydx)$ which have vanishing $y$–coordinates.
-The Legendrian arcs are allowed to enter the handles both from the left and right in the front projection, where the arcs are required to have constant $z$–coordinate.
+Our Kirby diagrams will consist of a number of attaching 0-spheres inside $(\mathbb{R}^{3}_{xyz},dz-ydx)$ which have vanishing yy–coordinates.
+The Legendrian arcs are allowed to enter the handles both from the left and right in the front projection, where the arcs are required to have constant zz–coordinate.
 The attaching 1-spheres will typically be depicted in red colour, while Legendrian knots will be black.
@@
 Move K–0: Legendrian isotopy in the complement of the attaching spheres: Both Legendrian knots in the contact manifold, as well as attaching 1-spheres, can be deformed by a Legendrian isotopy that supported in the complement of the attaching 0-spheres.
-By our convention the attaching 0-spheres have a vanishing $y$–coordinate, which means that any Legendrian arc whose front projection has non-zero slope can pass over the small balls where the attaching 0-spheres live.
+By our convention the attaching 0-spheres have a vanishing yy–coordinate, which means that any Legendrian arc whose front projection has non-zero slope can pass over the small balls where the attaching 0-spheres live.
 See Figure 3.
@@
 Move K–I: Introducing/cancelling a pair of handles: A Weinstein one- and two-handle that are in cancelling position can be either introduced or removed.
-This is a standard Kirby move; see [13, Section 5] in work by Ding and Geiges for the version in contact topology.
+This is a standard Kirby move; see [14, Section 5] in work by Ding and Geiges for the version in contact topology.
 This move is shown in Figure 4.
 Figure 4.
-The Kirby move K–I: The Kirby move that consists of cancelling a pair consisting of a Weinstein 1-handle and a Weinstein 2-handle in standard position corresponds to the following deformation of the front projection.
-Move K–II: Handle-sliding parallel strands: Handling sliding parallel strands of a Legendrian over the two-handle which is in cancelling position gives rise to a move shown in Figure 5.
+The Kirby move K–I: The Kirby move that consists of cancelling a pair consisting of a Weinstein 1-handle and a Weinstein 2-handle in standard position corresponds to the above deformation of the front projection.
+Move K–II: Handle-sliding parallel strands: Handling sliding parallel strands of a Legendrian over a Weinstein two-handle which is in cancelling position gives rise to a move shown in Figure 5.
 This is a standard Kirby move; the version in contact manifolds was described by Ding and Geiges in [14]; and Casals and Murphy in [2].
 Figure 5.
-The Kirby move K–II: The bottom figure shows the effect in the front projection of the Kirby move that consists of handle-sliding $k=2$ parallel Legendrian strands shown on top over its Legendrian two-handle whose attaching sphere is shown in red.
-Move K–III: Introducing a double-stabilisation: Taken any number of parallel strands that pass through a single one-handle, one can move the top one to the bottom, introducing a double stabilisation of the same.
-See [27, Figure 17] for one version, also its reflection in the z-axis is a possible move.
+The Kirby move K–II: The bottom figure shows the effect in the front projection of the Kirby move that consists of handle-sliding $k=2$ parallel Legendrian strands shown on top over the Weinstein two-handle whose attaching sphere is the knot shown in red.
+Move K–III: Introducing a double-stabilisation: Given any number of parallel strands that pass through a single one-handle, one can move the bottom strand to the top, introducing a double stabilisation of the same.
+See [28, Figure 17] for one version, also its reflection in the z-axis is a possible move.
 See Figure 6.
@@
 The strand is given an additional downward cusp-edge in the direction oriented away from the attaching sphere.
-See Figure 7.
+See Figure 7 for the move in the front projection, and Figure 8 for how the move can be described using the Lagrangian projection.
 Figure 7. The Kirby move K–IV: moving a strand entering the one handle from one side of the attaching sphere to the opposite side.
@@
 3.1. Decomposable fillings
-A particularly nice class of Lagrangian cobordisms in ×$\mathbb{R}\times S^{3}$ are the decomposable ones, which were defined by Chantraine in [4].
+A particularly nice class of Lagrangian cobordisms in $\mathbb{R}\times S^{3}$ are the decomposable ones, which were defined by Chantraine in [4].
 By definition, a Lagrangian cobordism is decomposable if it consists of a concatenation of the following simple types of Lagrangian cobordisms:
@@
 Figure 9.
-Ambient Legendrian surgery: The Legendrian is deformed in a neighborhood of the surgery arc (shown on top) to yield the Legendrian $\Lambda_{\eta}$ on the bottom.
-There is a standard Lagrangian 1-handle cobordism with concave boundary the Legendrian $\Lambda$ before the surgery, and convex boundary the Legendrian $\Lambda_{\eta}$ produced by the surgery.
+Ambient Legendrian surgery: The Legendrian $\Lambda$ is deformed in a neighborhood of the surgery arc ($\eta$ shown at the bottom) to yield the Legendrian $\Lambda_{+}=\Lambda_{\eta}$ shown on top.
+There is a standard Lagrangian 1-handle cobordism with concave boundary the Legendrian $\Lambda$ before the surgery, and convex boundary the Legendrian $\Lambda_{+}=\Lambda_{\eta}$ produced by the surgery.
 […]
@@
 cancelling the handle after the surgery by the K-II move again yields the Legendrian $\Lambda_{\eta}$ resulting by the surgery shown in the bottom of Figure 9.
-Decomposable Lagrangian cobordisms where shown to be regular in the work of Conway, Etnyre and Tosun [6].
+Decomposable Lagrangian cobordisms were shown to be regular in the work of Conway, Etnyre and Tosun [6].
 It is still unknown whether the converse is true or not.
@@
 We proceed to explain this recipe in the case of a decomposable Lagrangian filling.
-The 0-handles: For each zero-handle we consider $(\mathbb{C}^{2},-d^{c}\rho)$ with the standard filling $\mathfrak{Re}\mathbb{C}^{2}\subset\mathbb{C}^{2}$ of the standard Legendrian unknot $\partial_{\infty}\mathfrak{Re}\mathbb{C}^{2}\subset S^{3}$ of $\tt{tb}=-1$.
-This filling is clearly is tangent to the radial Liouville vector field, where $\rho(\mathbf{z})=\frac{\|\mathbf{z}\|^{2}}{4}.$ In the general case, when the Lagrangian cobordism has several 0-handles, one has to consider a Weinstein connected sum of such Weinstein four-balls.
+The 0-handles: For each zero-handle we consider $(\mathbb{C}^{2},-d^{c}\rho)$, where $\rho(\mathbf{z})=\frac{\|\mathbf{z}\|^{2}}{4},$ with the standard filling $\mathfrak{Re}\mathbb{C}^{2}\subset\mathbb{C}^{2}$ of the standard Legendrian unknot $\partial_{\infty}\left(\mathfrak{Re}\mathbb{C}^{2}\right)\subset S^{3}$ of $\tt{tb}=-1$.
+This filling is clearly is tangent to the radial Liouville vector field.
+In the general case, when the Lagrangian cobordism has several 0-handles, one has to consider a Weinstein connected sum of such Weinstein four-balls.
 Since this Weinstein connected sum again produces the standard sphere, and since these handles can be determined from the number of zero-handles, we will omit these 1-handles from the diagram.
@@
 The resulting diagram after having performed the cusp-connected sum is shown at the top of Figure 5.
-Before the cusp-connected sum, the two facing cusp-edges are now located inside the Weinstein 1-handle.
-In other words, the cusp-connected sum now takes place entirely inside the one-handle.
-It can readily be seen that the corresponding standard Lagrangian one-handle in this position is tangent to the standard Liouville structure on the Weinstein one-handle.
+Performing the K–II move as shown in the same figure places the two strands of the Legendrian that underwent cusp-connected sum inside the Weinstein 1-handle.
+(Alternatively, one could handle-slide the two facing cusp-edges together with the surgery arc into the one-handle similarly to the K–II move as shown in Figure 10.) It can readily be seen that the corresponding standard Lagrangian one-handle in this position can be made tangent to the standard Liouville structure on the Weinstein one-handle.
 In particular, the critical point of the Weinstein one-handle coincides with the critical point of the standard Lagrangian one-handle.
@@
 $\{0\}\times 0_{\eta}\times\{0\}\subset\{0\}\times 0_{\eta}\times 0_{D^{n-k-1}}\subset[-\epsilon,\epsilon]_{z}\times D^{*}\eta\times D^{*}D^{n-k-1}$
-of the zero-section in the above standard neighbourhood, while the Legendrian $\Lambda$ is identified with the Lagrangian co-normal bundle $N^{*}_{(\partial\eta)\times D^{n-k-1}}=N^{*}_{\partial\eta}\times 0_{D^{n-k-1}}$ of $(\partial\eta)\times D^{n-k-1}$, lifted to have vanishing $[-\epsilon,\epsilon]_{z}$-coordinate.
-The choice of conformal symplectic normal bundle induces the identification of the $D^{*}D^{n-k-1}$-factor in the above neighbourhood.
+of the zero-section in the above standard neighbourhood, while the Legendrian $\Lambda$ is identified with the Lagrangian co-normal bundle $N^{*}_{(\partial\eta)\times D^{n-k-1}}=N^{*}_{\partial\eta}\times 0_{D^{n-k-1}}$ of $(\partial\eta)\times D^{n-k-1}$, lifted to have vanishing $\epsilon,\epsilon]_{z}-$coordinate.
+The choice of conformal symplectic normal bundle induces an identification of the $D^{*}D^{n-k-1}$-factor in the above neighbourhood.
 Since $N^{*}_{(\partial\eta)\times D^{n-k-1}}$ is a trivial $\mathbb{R}$-fibration over $(\partial\eta)\times D^{n-k-1}$, the compatibility of the framing of $\partial\eta$ and the above conformal symplectic normal bundle implies that the $\mathbb{R}$-fibre of the conormal bundle together with the standard coordinates on the $0_{D^{n-k-1}}$-factor together yield the framing of $\partial\eta\subset\Lambda$.
-Passing to the symplectisation, we can find a neighbourhood of $[0,1]\times\eta\subset\mathbb{R}\times Y$ in the symplectisation of the form
+[…]
+[…]
+[…]
+Figure 11.
+The top depicts the front projection of the Legendrian link $\Lambda$ and the surgery arc $\eta$ in the standard jet-neighborhood of a surgery-arc.
+The middle shows the Legendrian $\Lambda^{\prime}$ obtained by a Legendrian isotopy of $\Lambda$ supported near $\eta$ that moves the two cusp-edges closer to each other.
+The bottom picture depicts the result of a cusp-connected sum of $\Lambda^{\prime}$.
+This process is the definition of an ambient Legendrian surgery $\Lambda_{\eta}$ of $\Lambda$ defined by the surgery arc $\eta$.
+Passing to the symplectisation, we can find a neighbourhood of $0,1]\times\eta\subset\mathbb{R}\times Y$ in the symplectisation of the form
 $D^{*}[0,1]\times D^{*}\eta\times D^{*}D^{n-k-1}$
-in which the cone $[0,1]\times\eta$ is identified with $0_{[0,1]}\times 0_{\eta}\times\{0\}$ and the cone $[0,1]\times\Lambda$ is identified with $0_{[0,1]}\times N^{*}_{\partial\eta}\times 0_{D^{n-k-1}}$.
+in which the cone $0,1]\times\eta$ is identified with $0_{[0,1]}\times 0_{\eta}\times\{0\}$ and the cone $0,1]\times\Lambda$ is identified with $0_{[0,1]}\times N^{*}_{\partial\eta}\times 0_{D^{n-k-1}}$.
 The standard Lagrangian $k+1$-handle is constructed to be a cylinder over $\Lambda$ outside the above neighbourhood of the symplectisation, while the intersection with the above neighbourhood is an explicitly constructed neighbourhood of the core-disc; see [9, Section 4.2.2].
@@
 $0_{[0,1]}\times 0_{\eta}\times 0_{D^{n-k-1}}\subset D^{*}[0,1]\times D^{*}\eta\times D^{*}D^{n-k-1}.$
-Namely, in a small neighbourhood of this zero-section, the standard handle $L_{\eta}$ is given by the Lagrangian co-normal bundle of a $k+1$-dimensional disc in $\Gamma\times\{0\}\subset[0,1]\times\eta\times D^{n-k-1}$ where $\Gamma\subset[0,1]\times\eta$ is a smooth codimension one embedding of a disc given as the union of $[0,1/2]\times\partial\eta$ and a suitable graph of a function $\eta\to[1/2,1]$.
+Namely, in a small neighbourhood of this zero-section, the standard handle $L_{\eta}$ is given by the Lagrangian co-normal bundle of a $k+1$-dimensional disc
+$\overline{C}_{\eta}=\Gamma\times\{0\}\subset[0,1]\times\eta\times D^{n-k-1}$
+where $\Gamma\subset[0,1]\times\eta$ is a smooth codimension one embedding of a disc given as the union of $0,1/2]\times\partial\eta$ and a suitable graph of a function $\eta\to[1/2,1]$.
 In particular, this means that $L_{\eta}$ intersects the zero-section of this standard neighbourhood in a representative $\overline{C}_{\eta}\subset\overline{L}_{\eta}$ of the $k+1$-dimensional core-disc of the handle.
@@
 The following lemma is a direct consequence of the construction in terms of a local model implemented in a choice of standard neighbourhood.
-Lemma 3.1.
+Lemma 3.1 ([9]).
 The Legendrian isotopy class of the resulting Legendrian $\Lambda_{\eta}$ only depends on $\Lambda\cup\eta$, together with the data of the framings, up to contact isotopy.
@@
   2. (2)
-    $L_{\eta}$ is contained in the $k+1$-handle and tangent to the Liouville flow, where the critical point of the Liouville flow of the handle coincides with the critical point of the restriction $t|_{L_{\eta}}$ of the symplectisation coordinate;
+    $L_{\eta}$ is contained in the Weinstein $k+1$-handle and is tangent to the Liouville flow, where the critical point of the Liouville flow of the handle coincides with the critical point of the restriction $t|_{L_{\eta}}$ of the symplectisation coordinate;
   3. (3)
@@
 the skeleton of the handle, also known as core-disc) is the isotropic $i$-disc $0_{D^{i}}\times\{0\}$, while its ascending manifold (i.e.
-co-core disc) is the coisotropic $2n-i$-disc $D^{*}_{0}D^{i}\times D^{2(n+1-i)}$.
-Moreover, the Liouville vector-field can be taken to be the radial Liouville vector field corresponding to the Liouville form $-d^{c}\|\mathbf{z}\|^{2}/4$ in the second component.
-We begin by symplectically identifying a neighbourhood of the isotropic $k+1$-disc $C_{\eta}$ with a standard Weinstein $k+1$-handle embedded in ×$\mathbb{R}\times Y$, which has been attached along the isotropic sphere $\partial\eta\subset\partial((-\infty,0]\times Y)$.
+co-core disc) is the coisotropic 2ni2n-i-disc $D^{*}_{0}D^{i}\times D^{2(n+1-i)}$.
+Moreover, the Liouville vector-field can be taken to be the radial Liouville vector field corresponding to the Liouville form $d^{c}\|\mathbf{z}\|^{2}/4$ in the second component.
+We begin by symplectically identifying a neighbourhood of the isotropic $k+1$-disc $C_{\eta}$ with a standard Weinstein $k+1$-handle embedded in $\mathbb{R}\times Y$, which has been attached along the isotropic sphere $\partial\eta\subset\partial((-\infty,0]\times Y)$.
 Here the framing of $\partial\eta$ in $Y$ is induced from the framing of $\partial\eta\subset\Lambda$.
@@
 (Note that this is different from the original Liouville vector field $\partial_{t}$ on the symplectisation.)
-Further, we may assume that this identification takes $L_{\eta}\cap\overline{X}_{a}$ to the Lagrangian
-$L_{0}\coloneqq 0_{D^{k+1}}\times(\mathfrak{Im}\mathbb{C}^{n-k}\cap D^{2(n-k)})\subset(D^{*}D^{k+1}\times D^{2(n-k)},d(p\,dq)+\omega_{0})$
-contained in the standard Weinstein $k+1$-handle that is everywhere tangent to the Liouville vector field, while $D_{\eta}\cap\overline{X}_{a}$ inside the handle is identified with
-$0_{D^{k+1}}\times\left(\{x_{1}\geq 0,x_{2}=\ldots=x_{n-k}=y_{1}=\ldots=y_{n-k}=0\}\cap D^{2(n-k)}\right).$
-The intersection of $D_{\eta}$ with the contact boundary $\partial\overline{X}_{a}\subset\mathbb{R}\times Y$ can, moreover, be assumed to be an isotropic $k+1$-sphere.
-We may symplectically identify a neighbourhood of $D_{\eta}\setminus X_{a}$ with a Weinstein $k+2$-handle, under which $D_{\eta}\setminus C_{\eta}$ is the isotropic core-disc of the handle.
-Using $\overline{X}_{b}\supset\overline{X}_{a}$ to denote the union of $\overline{X}_{a}$ together with a neighbourhood of $D_{\eta}$, we have produced a symplectic identification of the neighbourhood $\overline{X}_{b}\subset\mathbb{R}\times Y$ with the result of attaching Weinstein handles of indices $k+1$ and $k+2$ on $\partial(-\infty,0]\times Y$ in cancelling position.
-In particular, $\partial\overline{X}_{b}$ is contactomorphic to $Y$.
-Finally, after a compactly supported deformation of the Liouville vector field $\partial_{t}$ of (×$(\mathbb{R}\times Y,d(e^{t}\alpha))$ to a complete Liouville vector field $\zeta$ without critical points for the same symplectic form, we may assume that $L_{\eta}\setminus X_{b}$ is tangent to $\zeta$ in ×$\mathbb{R}\times Y\setminus\overline{X}_{b}$, and that $\zeta$ coincides with the Liouville vector field induced by the Weinstein handle-decomposition of $\overline{X}_{b}$ near its boundary $\partial\overline{X}_{b}$.
-This gives us the sought Weinstein structure on ×$\mathbb{R}\times Y$ and thus finishes the proof of properties (1)–(4).
-In the special case when $n=1$ and $k=0$, the handles in cancelling position together with the Legendrian $\Lambda_{\eta}$ consists of:
-  • A Weinstein one-handle attached at the 0-sphere $\partial\eta$;
-  • A Weinstein two-handle attached along a knot $S_{\eta}$ that coincides with $\eta$ outside of the above 00-handle; while
-  • The Legendrian $\Lambda_{\eta}$ coincides with $\Lambda$ outside of the one-handle and coincides with a cusp-connected sum of two strands inside the one-handle.
-In Darboux balls that contain $\partial\eta$ the Legendrian $\Lambda_{\eta}$ and attaching 00-spheres and $1$-spheres $S_{\eta}$ can be described as shown on the bottom in Figure 13.
-Note that this Kirby diagram is not in standard form, since there are strands entering the one-handle from both the left and right.
-This we correct by the contact isotopy described in K–IV move applied to $\Lambda_{\eta}$; see Figure 7.
-This contact isotopy is supported inside the one-handle, fixes the attaching sphere $S_{\eta}$ of the cancelling two-handle, and places the strands that correspond to $\Lambda_{\eta}$ on the same side as the strand corresponding to $S_{\eta}$.
-This produces the configuration shown on the bottom of Figure 14, as sought.
-∎
-3.3. Ambient surgery, cusp-connected sum, and isotopies.
-Since the main focus here is Legendrian ambient 00-surgery on Legendrian knots in $\mathbb{R}^{3}$ we will give some additional details in this case.
-When $\eta$ connects two cusp edges in the front projection of the Legendrian link $\Lambda$, the Legendrian ambient surgery can be performed in the following manner.
-Step (1): Choose a standard contact jet-neighbourhood $J^{1}\eta$ of the surgery arc in which the Legendrian $\Lambda$ consists of two facing cusp-edges, whose singularities are connected by the zero-section $\eta\subset J^{1}\eta$ that corresponds to the arc $\eta$.
-Step (2): Perform a Legendrian isotopy of $\Lambda$ whose support is compact inside the above neighbourhood $J^{1}\eta$, so that the cusp-edges become arbitrarily close to each other; denote the resulting Legendrian by $\Lambda^{\prime}$.
-We perform this isotopy in an arbitrarily small neighborhood of $\eta$, while fixing the arc $\eta$ set-wise, and so that it induces an isotopy of front-diagrams; See the top and middle front in Figure 11.
-Step (3): Perform a standard cusp-connected sum at the two facing cusp-edges of $\Lambda^{\prime}$ that now are joined by a very small segment of $\eta$, and can be assumed to be arbitrarily close.
-The result is shown in the bottom of Figure 11.
-[…]
-[…]
-[…]
-Figure 11.
-The top depicts the front projection of the Legendrian link $\Lambda$ and the surgery arc $\eta$ in the standard jet-neighborhood of a surgery-arc.
-The middle shows the Legendrian $\Lambda^{\prime}$ obtained by a Legendrian isotopy of $\Lambda$ supported near $\eta$ that moves the two cusp-edges closer to each other.
-The bottom picture depicts the result of a cusp-connected sum of $\Lambda^{\prime}$.
-This process is the definition of an ambient Legendrian surgery $\Lambda_{\eta}$ of $\Lambda$ defined by the surgery arc $\eta$.
-[…]
-[…]
-[…]
-Figure 12. The Legendrian isotopy of $\Lambda$ following the surgery arc
-Moreover, the ambient Legendrian surgery gives rise to a Lagrangian standard handle-attachment cobordism that consists of the trace of the aforementioned isotopy, followed by a Lagrangian standard one-handle.
-The isotopy performed in Step (2) is easy to describe also in the front of $\mathbb{R}^{3}$; we express this in a lemma.
-Lemma 3.3.
-For a generic choice of surgery arc $\eta$ and Legendrian $\Lambda$, the result $\Lambda_{\eta}$ after Legendrian ambient surgery can be obtained by first applying the Legendrian isotopy of $\Lambda$ in Step (2) by performing consecutive moves of type (S1) and (S2) shown in Figure 12 to yield $\Lambda^{\prime}$, and then performing a standard cusp-connected sum on the resulting $\Lambda^{\prime}$.
 […]
@@
 Figure 13.
-The Kirby diagram on the bottom is obtained from the one on the bottom of Figure 13 by a contact isotopy which is the identity in a neighbourhood of $S_{\eta}$, is supported in the one-handle, and which moves $\Lambda_{\eta}$ so that it enters the one-handle on the same side as $S_{\eta}$, with the two strands having larger $z$–coordinates.
+The Kirby diagram on the bottom is obtained from the one on the bottom of Figure 12 by a contact isotopy which is the identity in a neighbourhood of $S_{\eta}$, is supported in the one-handle, and which moves $\Lambda_{\eta}$ so that it enters the one-handle on the same side as $S_{\eta}$, with the two strands having larger zz–coordinates.
+Further, we may assume that this identification takes $L_{\eta}\cap\overline{X}_{a}$ to the Lagrangian
+$L_{0}\coloneqq 0_{D^{k+1}}\times(\mathfrak{Im}\mathbb{C}^{n-k}\cap D^{2(n-k)})\subset(D^{*}D^{k+1}\times D^{2(n-k)},d(p\,dq)+\omega_{0})$
+contained in the standard Weinstein $k+1$-handle that is everywhere tangent to the Liouville vector field, while $D_{\eta}\cap\overline{X}_{a}$ inside the handle is identified with
+$0_{D^{k+1}}\times\left(\{x_{1}\geq 0,x_{2}=\ldots=x_{n-k}=y_{1}=\ldots=y_{n-k}=0\}\cap D^{2(n-k)}\right).$
+The intersection of $D_{\eta}$ with the contact boundary $\partial\overline{X}_{a}\subset\mathbb{R}\times Y$ can, moreover, be assumed to be an isotropic $k+1$-sphere.
+We may symplectically identify a neighbourhood of $D_{\eta}\setminus X_{a}$ with a Weinstein $k+2$-handle, under which $D_{\eta}\setminus C_{\eta}$ is the isotropic core-disc of the handle.
+Using $\overline{X}_{b}\supset\overline{X}_{a}$ to denote the union of $\overline{X}_{a}$ together with a neighbourhood of $D_{\eta}$, we have produced a symplectic identification of the neighbourhood $\overline{X}_{b}\subset\mathbb{R}\times Y$ with the result of attaching Weinstein handles of indices $k+1$ and $k+2$ on $\partial(-\infty,0]\times Y$ in cancelling position.
+In particular, $\partial\overline{X}_{b}$ is contactomorphic to $Y$.
+Finally, after a compactly supported deformation of the Liouville vector field $\partial_{t}$ of $(\mathbb{R}\times Y,d(e^{t}\alpha))$ to a complete Liouville vector field $\zeta$ without critical points for the same symplectic form, we may assume that $L_{\eta}\setminus X_{b}$ is tangent to $\zeta$ in $\mathbb{R}\times Y\setminus\overline{X}_{b}$, and that $\zeta$ coincides with the Liouville vector field induced by the Weinstein handle-decomposition of $\overline{X}_{b}$ near its boundary $\partial\overline{X}_{b}$.
+This gives us the sought Weinstein structure on $\mathbb{R}\times Y$ and thus finishes the proof of properties (1)–(4).
+In the special case when $n=1$ and $k=0$, the handles in cancelling position together with the Legendrian $\Lambda_{\eta}$ consists of:
+  • A Weinstein one-handle attached at the 0-sphere $\partial\eta$;
+  • A Weinstein two-handle attached along a knot $S_{\eta}$ that coincides with $\eta$ outside of the above 0-handle; while
+  • The Legendrian $\Lambda_{\eta}$ coincides with $\Lambda$ outside of the one-handle and coincides with a cusp-connected sum of two strands inside the one-handle.
+In Darboux balls that contain $\partial\eta$ the Legendrian $\Lambda_{\eta}$ and attaching 0-spheres and $1$-spheres $S_{\eta}$ can be described as shown on the bottom in Figure 12.
+Note that this Kirby diagram is not in standard form, since there are strands entering the one-handle from both the left and right.
+This we correct by the contact isotopy described in K–IV move applied to $\Lambda_{\eta}$; see Figure 7.
+This contact isotopy is supported inside the one-handle, fixes the attaching sphere $S_{\eta}$ of the cancelling two-handle, and places the strands that correspond to $\Lambda_{\eta}$ on the same side as the strand corresponding to $S_{\eta}$.
+This produces the configuration shown on the bottom of Figure 13, as sought.
+∎
+3.3. Ambient surgery, cusp-connected sum, and isotopies.
+Since the main focus here is Legendrian ambient 0-surgery on Legendrian knots in $\mathbb{R}^{3}$ we will give some additional details in this case.
+When $\eta$ connects two cusp edges in the front projection of the Legendrian link $\Lambda$, the Legendrian ambient surgery can be performed in the following manner.
+Step (1): Choose a standard contact jet-neighbourhood $J^{1}\eta$ of the surgery arc in which the Legendrian $\Lambda$ consists of two facing cusp-edges, whose singularities are connected by the zero-section $\eta\subset J^{1}\eta$ that corresponds to the arc $\eta$.
+Step (2): Perform a Legendrian isotopy of $\Lambda$ whose support is compact inside the above neighbourhood $J^{1}\eta$, so that the cusp-edges become arbitrarily close to each other; denote the resulting Legendrian by $\Lambda^{\prime}$.
+We perform this isotopy in an arbitrarily small neighborhood of $\eta$, while fixing the arc $\eta$ set-wise, and so that it induces an isotopy of front-diagrams; See the top and middle front in Figure 11.
+Step (3): Perform a standard cusp-connected sum at the two facing cusp-edges of $\Lambda^{\prime}$ that now are joined by a very small segment of $\eta$, and can be assumed to be arbitrarily close.
+The result is shown in the bottom of Figure 11.
+Moreover, the ambient Legendrian surgery gives rise to a Lagrangian standard handle-attachment cobordism that consists of the trace of the aforementioned isotopy, followed by a Lagrangian standard one-handle.
+The isotopy performed in Step (2) is easy to describe also in the front of $\mathbb{R}^{3}$; we express this in a lemma.
+The techniques, and moves shown in the figures, are covered by the generalised Reidemeister moves for Legendrian graphs developed in [35], which also have been used by Sabloff, Vela-Vick, Wong, and Wu in [38].
+Lemma 3.3.
+For a generic choice of surgery arc $\eta$ and Legendrian $\Lambda$, the result $\Lambda_{\eta}$ after Legendrian ambient surgery can be obtained by first applying the Legendrian isotopy of $\Lambda$ in Step (2) by performing consecutive moves of type (S1) and (S2) shown in Figure 14 to yield $\Lambda^{\prime}$, and then performing a standard cusp-connected sum on the resulting $\Lambda^{\prime}$.
+[…]
+[…]
+[…]
+Figure 14. The Legendrian isotopy of $\Lambda$ following the surgery arc. These moves were also established by Sabloff, Vela-Vick, Wong and Wu, see e.g. [38, Figure 9] for (S2).
 4. Spinning surgery discs and attaching spheres (Proof of Theorem 1.3)
@@
 These strategies will lead to the proof of Theorem 1.3. We will mainly be interested in the case then $Y^{2m+1}=\mathbb{R}^{3}$. We begin by showing some results that can be of general use.
-Lemma 4.1.
-Assume that $(Y,\xi=\ker\alpha)$ is a contact manifold for which any proper compact subset $K\subsetneq Y$ admits a displacing contact isotopy $\phi_{t}\in\mathrm{Cont}(Y,\xi)$, by which we mean that $\phi_{1}(K)\cap K=\emptyset$.
-(This is e.g.
-the case for $Y=\mathbb{R}^{2m+1}$ or the standard contact $S^{2m+1}$.) Then, if $\Lambda\subset(Y\times T^{*}S^{n},\alpha+p\,d\theta)$ is a closed Legendrian, the image $\Psi(\Lambda)$ is Legendrian isotopic to $\Lambda$ whenever $\Psi\in\mathrm{Cont}(Y\times T^{*}S^{n})$ is a contactomorphism with support contained in $K\times T^{*}S^{n}$ for some proper compact subset $K\subsetneq Y$.
-Proof.
-Any contact isotopy $\phi_{t}\in\mathrm{Cont}(Y,\xi)$ lifts to a contact isotopy
-$\Phi_{t}\in\mathrm{Cont}(Y\times T^{*}S^{n}),$
-$\Phi_{t}(y,\theta,p)=(\phi_{t}(y),\theta,p\cdot e^{g_{t}(y)})$
-where $g_{t}\colon Y\to\mathbb{R}$ is determined by $\phi_{t}^{*}\alpha=e^{g_{t}}\alpha$.
-For any contactomorphism $\Psi\in\mathrm{Cont}(Y\times T^{*}S^{n})$ supported in $K\times T^{*}S^{n}$ with $K\subsetneq Y$ proper and compact, we can find a contact isotopy of the above form for which $\Phi_{1}$ displaces the support of $\Psi$.
-The image $\Psi(\Lambda)$ is thus Legendrian isotopic to $\Psi\circ\Phi_{1}(\Lambda)=\Phi_{1}(\Lambda)$ which, in turn, is Legendrian isotopic to $\Phi_{0}(\Lambda)=\Lambda$.
-∎
+4.1. Generalities on formal Legendrian isotopies in the complement of a Legendrian
 Proposition 4.1.
-Assume that two Legendrian links (resp.
-union of properly embedded arcs) $S_{0},S_{1}\subset Y^{3}\setminus\Lambda$ admit a (compactly supported) Legendrian isotopy $S_{t}\subset Y$, where $\Lambda\subset Y$ is a Legendrian link.
-It follows that:
+Let $\Lambda\subset Y^{3}$ denote a fixed Legendrian link.
+Further, let $S_{0},S_{1}\subset Y^{3}\setminus\Lambda$ be two Legendrians in the complement of $\Lambda$, where each component of $S_{i}$ is either closed or compact with boundary.
+If there exists a Legendrian isotopy $S_{t}\subset Y$ which takes $S_{0}$ to $S_{1}$ (the isotopy is allowed to cross $\Lambda$), which is the identity near the boundary $\partial S_{0}\cup\partial S_{1}$, then it follows that:
   • When $n=2l+1>0$ is odd, the two products $S_{i}\times 0_{S^{n}}$, $i=0,1$, are (compactly supported) formally Legendrian isotopic when considered inside the contact manifold $(Y\times T^{*}S^{n})\setminus(\Lambda\times 0_{S^{n}})$.
@@
     $\{(t,y);\>\>y\in S_{t}\cup\Lambda\}\subset[0,1]\times Y$
-    has vanishing algebraic intersection number.
+    have vanishing algebraic intersection number.
 Proof.
@@
 We then deform the isotopy $S_{t}\times 0_{S^{n}}$ through isotopies to one that avoids $\Lambda\times 0_{S}^{n}$ by the following argument using the existence of certain Whitney discs:
-Create Whitney disc by connecting two double points of opposite signs with arcs in both intersecting sheets of the immersed trace cobordism $\mathcal{T}$; this forms a loop in $[0,1]_{t}\times Y\times T^{*}S^{n}$.
+Create Whitney disc by connecting two double points of opposite signs with arcs in both intersecting sheets of the immersed trace cobordism $\mathcal{T}$; this forms a loop in $0,1]_{t}\times Y\times T^{*}S^{n}$.
 We may assume that each arc intersects each constant $t$-slice $\{t\}\times Y\times T^{*}S^{n}$ in at most one point.
@@
 ∎
-Using the fact that products of stabilised knots become loose [10, Proposition 1.13], combined with Murphy’s $h$-principle for loose Legendrian embeddings [32], we get:
+Using the fact that products of stabilised knots become loose [11, Proposition 1.13], combined with Murphy’s $h$-principle for loose Legendrian embeddings [34], we get:
 Corollary 4.2.
@@
 Corollary 4.3.
-If the two Legendrians $S_{0},S_{1}\subset Y^{3}\setminus\Lambda$ satisfy the assumptions of Proposition 4.2 and, moreover, are stabilised when considered inside the contact three-manifold $Y^{3}\setminus\Lambda$, then the two Legendrians $\Lambda\times 0_{S^{n}}\subset Y_{S_{i}}\times T^{*}S^{n}$, $i=0,1$, are contactomorphic.
+If the two Legendrians $S_{0},S_{1}\subset Y^{3}\setminus\Lambda$ satisfy the assumptions of Proposition 4.1 and, moreover, are stabilised when considered inside the contact three-manifold $Y^{3}\setminus\Lambda$, then the two Legendrians $\Lambda\times 0_{S^{n}}\subset Y_{S_{i}}\times T^{*}S^{n}$, $i=0,1$, are compactly supported contactomorphic.
 When $Y_{S_{i}}\cong\mathbb{R}^{3}$, then the induced Legendrians inside $\mathbb{R}^{3}\times T^{*}S^{n}$ are Legendrian isotopic.
+The existence of a Legendrian isotopy in $\mathbb{R}^{3}\times T^{*}S^{n}$ is a direct consequence of the existence of a compactly supported contact contactomorphism.
+Namely, since any compact subset in $\mathbb{R}^{3}$ can be displaced from itself by a contactomorphism, any compactly supported contactomorphism of $\mathbb{R}^{3}\times T^{*}M$ preserves the Legendrian isotopy class of any compact Legendrian; see Lemma 4.5 below.
 We are now ready to prove Theorem 1.3, which follows from either one of the following two constructions.
@@
 Since standard neighbourhoods are unique up to contact isotopy, the construction does not depend on the choice of neighbourhood.
-Since there is a canonical identification $(J^{1}\eta_{i}\times T^{*}S^{n},dz+p\,dq+\lambda_{S^{n}})=J^{1}(\eta_{i}\times S^{n}),dz+p\,dq+\lambda_{S^{n}})$ of contact manifolds, a product of a contact standard neighbourhood of $\eta_{i}$ with a Weinstein neighbourhood of $0_{S}^{n}$ is a standard neighbourhood of the Legendrian $\eta_{i}\times 0_{S^{n}}$.
+Since there is a canonical identification
+$(J^{1}\eta_{i}\times T^{*}S^{n},dz+p\,dq+\lambda_{S^{n}})=J^{1}(\eta_{i}\times S^{n}),dz+p\,dq+\lambda_{S^{n}})$
+of contact manifolds, a product of a contact standard neighbourhood of $\eta_{i}$ with a Weinstein neighbourhood of $0_{S}^{n}$ is a standard neighbourhood of the Legendrian $\eta_{i}\times 0_{S^{n}}$.
 Any two contact neighbourhoods of $\eta_{i}\times 0_{S^{n}}$, $i=0,1,$ are contact isotopic; here we use the uniqueness of contact neighbourhood together with the assumption that these two Legendrians are isotopic.
@@
 In particular, the ambient Legendrian surgery produces the Legendrian knot $\Lambda_{\eta_{m}}=\Lambda_{m}$.
-Then consider the Legendrian surgery arc $\eta_{0}$ for the same Legendrian link $\Lambda$ depicted in Figure 19, for which $\Lambda_{\eta_{0}}$ clearly is Legendrian isotopic to the standard unknot $U$.
-Note that $\eta_{0}$ and $\eta_{m}$ both are stabilised and Legendrian isotopic in $\mathbb{R}^{3}$ relative boundary.
+Then consider the Legendrian surgery arc $\eta^{0}_{m}$ for the same Legendrian link $\Lambda$ depicted in Figure 19, for which $\Lambda_{\eta^{0}_{m}}$ clearly is Legendrian isotopic to the standard unknot $U$.
+Note that $\eta^{0}_{m}$ and $\eta_{m}$ both are stabilised and Legendrian isotopic in $\mathbb{R}^{3}$ relative boundary.
 However, any isotopy must clearly intersect $\Lambda$.
-In the case when $n$ is odd, we can apply Corollary 4.3 directly to get the sought isotopy between the spun of $\eta_{m}$ and the spun of $\eta_{0}$.
+In the case when $n$ is odd, we can apply Corollary 4.2 directly to get the sought isotopy between the spun of $\eta_{m}$ and the spun of $\eta^{0}_{m}$.
 Lemma 4.4 then gives the sought Legendrian isotopy between the spuns.
-In the case when $n$ is even, we can use the contact isotopy depicted in Figure 15 in order to verify the existence of a Legendrian isotopy that takes the arc $\eta_{m}$ to $\eta_{0}$, and for which the assumptions of Proposition 4.2 are met.
+In the case when $n$ is even, we can use the contact isotopy depicted in Figure 15 in order to verify the existence of a Legendrian isotopy that takes the arc $\eta_{m}$ to $\eta^{0}_{m}$, and for which the assumptions of Proposition 4.1 are met.
 Indeed, the obvious isotopy $\eta_{t}$ described in Figure 16 intersects $\Lambda$ twice, but with opposite signs of the corresponding double points on the induced trace cobordism.
-Again, corollary 4.3 finally shows the claim.
+Again, Corollary 4.2 shows the claim.
 […]
@@
 the result $\Lambda_{\eta_{m}}$ of an ambient Legendrian surgery on $\Lambda$ along the isotropic arc $\eta_{m}$.
-The result is a Weinstein structure on ×$\mathbb{R}\times\mathbb{R}^{3}$ consisting of a Weinstein one and two-handle in cancelling position for which the Lagrangian handle-attachment cobordism is tangent to the Liouville flow.
-In particular Proposition 3.2 gives us a Legendrian $\tilde{\Lambda}_{\eta}\subset\mathbb{R}^{3}_{sc}$ where the latter contact manifold is the result of a Weinstein 00-surgery on $\mathbb{R}^{3}$, such that $\tilde{\Lambda}_{\eta_{m}}$ becomes $\Lambda_{\eta_{m}}$ after a Weinstein 1-surgery on a knot $S\subset\mathbb{R}^{3}_{sc}\setminus\tilde{\Lambda}_{\eta_{m}}$ that cancels the 00-surgery.
-The attaching 00-sphere and cancelling $1$-sphere, as well as the knot $\tilde{\Lambda}_{\eta_{m}}$, are obtained from $\Lambda$ and $\eta_{m}$ by the recipe shown in Figure 14.
-There exists a stabilised Legendrian knot $S_{0}\subset(\mathbb{R}^{3})_{sc}\setminus\Lambda_{\eta_{m}}$ which also is in cancelling position, but for which $\Lambda_{\eta_{m}}$ becomes the standard unknot $U$ after cancelling the handles.
-For instance, we can use the ambient Legendrian surgery arc from Figure 19 and replace it with the corresponding attaching sphere from Proposition 3.2; again we can follow the recipe in Figure 14.
-Taking products with $T^{*}S^{n}$ results in a generalised Weinstein handle decomposition where $\mathbb{R}^{3}_{sc}\times T^{*}S^{n}$ are obtained from $\mathbb{R}^{3}\times T^{*}S^{n}$ by two subcritical surgeries along isotropic $n$-spheres, and where $\mathbb{R}^{3}\times T^{*}S^{n}$ is obtained from the latter by a generalised Weinstein surgery along the Legendrian $S\times 0_{S^{n}}$; this generalised surgery corresponds to attaching a generalised Weinstein handle of the form $T^{*}(D^{2}\times S^{n})$.
+The result is a Weinstein structure on $\mathbb{R}\times\mathbb{R}^{3}$ consisting of a Weinstein one and two-handle in cancelling position for which the Lagrangian handle-attachment cobordism is tangent to the Liouville flow.
+In particular Proposition 3.2 gives us a Legendrian $\tilde{\Lambda}_{\eta_{\color[rgb]{0,0,0}m}}\subset\left(\mathbb{R}^{3}\right)_{sc}$ where the latter contact manifold is the result of a Weinstein 0-surgery on $\mathbb{R}^{3}$, and such that $\tilde{\Lambda}_{\eta_{m}}$ becomes $\Lambda_{\eta_{m}}$ after a Weinstein 1-surgery on a knot $S\subset\left(\mathbb{R}^{3}\right)_{sc}\setminus\tilde{\Lambda}_{\eta_{m}}$ that cancels the 0-surgery.
+The attaching 0-sphere and cancelling $1$-sphere, as well as the knot $\tilde{\Lambda}_{\eta_{m}}$, are obtained from $\Lambda$ and $\eta_{m}$ by the recipe shown in Figure 13; we refer to Figure 18 for $\eta_{m}$.
+There exists a stabilised Legendrian knot $S_{0}\subset\left(\mathbb{R}^{3}\right)_{sc}\setminus\Lambda_{\eta_{m}}$ which also is in cancelling position, but for which $\Lambda_{\eta_{m}}$ becomes the standard unknot $U$ after cancelling the 1-handle with the 2-handle attached along $S_{0}$.
+This attaching sphere can be constructed analogously to $S$, but using the surgery arc $\eta^{0}_{m}$ shown in Figure 19 instead of $\eta_{m}$ shown in Figure 18.
+Here we again follow the recipe in Figure 13.
+4.3.1. Digresson: Generalised Weinstein handle decompositions
+The product of two Weinstein manifolds are naturally equipped with a handle-body structure where the handles of the product corresponds to products of handles.
+In particular, the number of handles for the induced handle-decomposition is the product of the number of handles of each factor.
+In the case when one of the factors is the cotangent bundle $T^{*}M$ of a closed smooth manifold there is a more economic decomposition, if one allows a generalisation of the Weinstein handle decomposition.
+First recall that the standard Weinstein $k$-handle, as introduced by Weinstein in [41], in dimension $2n$ with $n\geq k$ is given by $T^{*}D^{k}\times\mathbb{C}^{n-k}$ with the standard symplectic structure.
+This handle can be attached to a Liouville domain by, roughly speaking, gluing the k1k-1-sphere
+$\partial 0_{D^{k}}\times\{0\}\subset T^{*}D^{k}\times\mathbb{C}^{n-k}$
+to a framed isotropic k1k-1 sphere in the boundary of a Liouville domain.
+The latter framed isotropic embedding is called the isotropic attaching sphere.
+A Weinstein manifold can be characterised as a Liouville domain that admits a handle-body decomposition consisting of standard Weinstein handles.
+A generalised Weinstein $k$-handle is given by $T^{*}N^{k}\times\mathbb{C}^{n-k}$ where $D^{k}$ has been replaced by a general compact and smooth $k$-dimensional manifold $N^{k}$ with non-empty boundary.
+The generalised handles are similarly attached to a Liouville domain by gluing
+$\partial 0_{N}\times\{0\}\subset T^{*}M^{k}\times\mathbb{C}^{n-k}$
+along a framed isotropic embedding of $\partial N$ in the the contact boundary of the Liouville domain.
+We refer to the latter framed isotropic embedding as the isotropic attaching manifold; see [10] and [30] for more details on particular generalised handles.
+Those Liouville manifolds that admit a handle-body structure consisting of generalised Weinstein handles are obviously also Weinstien manifolds.
+By a generalised Weinstein handle body we mean a decomposition of a Weinstein manifold into generalised Weinstein handles.
+A particular set of examples that admit natural generalised Weinstein handle decompositions is the product $X\times T^{*}M^{m}$ of a Weinstein manifold $X$ and a cotangent bundle $T^{*}M$ of a closed smooth $m$-dimensional manifold.
+Namely, this product has a simple decomposition into generalised Weinstein handles induced by any Weinstein handle decomposition of $X$.
+Namely, each Weinstein handle $T^{*}D^{k}\times\mathbb{C}^{n-k}$ in the handle-decomposition of $X$ gives rise to a generalised Weinstein handle $T^{*}(D^{k}\times M)\times\mathbb{C}^{n-k}$ in the generalised Weinstein handle-body decomposition of $X\times T^{*}M$.
+The topology of each handle in this case is thus $N^{k+m}=D^{k}\times M^{m}$.
+4.3.2. A generalised Weinstein handle decomposition of the symplectisation $\mathbb{R}\times(\mathbb{R}^{3}\times T^{*}S^{n})$.
+In our situation we are interested in taking the product with $T^{*}S^{n}$.
+This results in a generalised Weinstein handle decomposition of the symplectisation $\mathbb{R}\times(\mathbb{R}^{3}\times T^{*}S^{n})$ obtained in the following manner.
+First, consider a Weinstein 1-handle attached on $\mathbb{R}^{3}$ and take the product with $T^{*}S^{n}$.
+This leads to a product Weinstein manifold with boundary $\left(\mathbb{R}^{3}\right)_{sc}\times T^{*}S^{n}.$ The latter contact manifold is obtained from $\mathbb{R}^{3}\times T^{*}S^{n}$ by a pair of subcritical Weinstein surgeries along isotropic $n$-spheres.
+Since the Weinstein 1-handle attachment that produces $\left(\mathbb{R}^{3}\right)_{sc}$ can be cancelled by a Weinstein 2-handle attached along $S\subset\left(\mathbb{R}^{3}\right)_{sc}$, taking a product of the latter handle and $T^{*}S^{n}$ then yields $\mathbb{R}^{3}\times T^{*}S^{n}$ obtained from $\left(\mathbb{R}^{3}\right)_{sc}\times T^{*}S^{n}$ by a generalised Weinstein surgery along the Legendrian $S\times 0_{S^{n}}$.
+Note that the latter generalised surgery corresponds to attaching a generalised Weinstein handle of the form $T^{*}(D^{2}\times S^{n})$.
 Note that such a handle-attachment, as well as the corresponding contact surgery, is completely determined by a choice of Legendrian embedding of $D^{2}\times S^{n}$.
-Finally we can apply Corollary 4.4 to the two attaching tori $S_{0}\times 0_{S^{n}}$ and $S\times 0_{S^{n}}$ to deduce the claim that $\Lambda_{m}\times 0_{S^{n}}$ and $U\times 0_{S^{n}}$ are Legendrian isotopic.
-Alternatively, one can invoke the $h$-principle for flexible Lagrangian cobordisms [18, Theorem 4.2] to deduce that there is a symplectomorphism of the symplectisation of $\mathbb{R}^{3}\times T^{*}S^{n}$ that identifies the standard Lagrangian fillings of $U\times 0_{S^{n}}$ with the one of $\Lambda_{m}\times 0_{S^{n}}$ (given by the product of the one in Figure 2).
-The Legendrian isotopy is then constructed by applying Lemma 4.1 to the contactomorphism of the contact boundary that is induced by the symplectomorphism.
+Finally we can apply Corollary 4.3 to the two attaching manifolds $S_{0}\times 0_{S^{n}}$ and $S\times 0_{S^{n}}$ to deduce the claim that $\Lambda_{m}\times 0_{S^{n}}$ and $U\times 0_{S^{n}}$ are Legendrian isotopic.
+Alternatively, one can invoke the $h$-principle for flexible Lagrangian cobordisms [19, Theorem 4.2] to deduce that there is a symplectomorphism of the symplectisation of $\mathbb{R}^{3}\times T^{*}S^{n}$ that identifies the standard Lagrangian fillings of $U\times 0_{S^{n}}$ with the one of $\Lambda_{m}\times 0_{S^{n}}$ (given by the product of the one in Figure 2).
+The Legendrian isotopy is then constructed by applying the below Lemma 4.5 to the contactomorphism of the contact boundary that is induced by the symplectomorphism.
+Lemma 4.5.
+Assume that $(Y,\xi=\ker\alpha)$ is a contact manifold for which any proper compact subset $K\subsetneq Y$ admits a displacing contact isotopy $\phi_{t}\in\mathrm{Cont}(Y,\xi)$, by which we mean that $\phi_{1}(K)\cap K=\emptyset$.
+(This is e.g.
+the case for $Y=\mathbb{R}^{2m+1}$ or the standard contact $S^{2m+1}$.) Then, if $\Lambda\subset(Y\times T^{*}S^{n},\alpha+p\,d\theta)$ is a closed Legendrian, the image $\Psi(\Lambda)$ is Legendrian isotopic to $\Lambda$ whenever $\Psi\in\mathrm{Cont}(Y\times T^{*}S^{n})$ is a contactomorphism with support contained in $K\times T^{*}S^{n}$ for some proper compact subset $K\subsetneq Y$.
+Proof.
+Any contact isotopy $\phi_{t}\in\mathrm{Cont}(Y,\xi)$ lifts to a contact isotopy
+$\Phi_{t}\in\mathrm{Cont}(Y\times T^{*}S^{n}),$
+$\Phi_{t}(y,\theta,p)=(\phi_{t}(y),\theta,p\cdot e^{g_{t}(y)})$
+where $g_{t}\colon Y\to\mathbb{R}$ is determined by $\phi_{t}^{*}\alpha=e^{g_{t}}\alpha$.
+For any contactomorphism $\Psi\in\mathrm{Cont}(Y\times T^{*}S^{n})$ supported in $K\times T^{*}S^{n}$ with $K\subsetneq Y$ proper and compact, we can find a contact isotopy of the above form for which $\Phi_{1}$ displaces the support of $\Psi$.
+The image $\Psi(\Lambda)$ is thus Legendrian isotopic to $\Psi\circ\Phi_{1}(\Lambda)=\Phi_{1}(\Lambda)$ which, in turn, is Legendrian isotopic to $\Phi_{0}(\Lambda)=\Lambda$.
+∎
 […]
@@
 Figure 19.
-Performing the Legendrian ambient surgery on two unlinked unknots, whose front projections moreover are disjoint, along an arc $\eta_{0}$ that intersects the two unknots precisely in their boundary points, produces the standard unknot $U$.
+Performing the Legendrian ambient surgery on two unlinked unknots, whose front projections moreover are disjoint, along an arc $\eta^{0}_{m}$ that intersects the two unknots precisely in their boundary points, produces the standard unknot $U$.
 5. Constructing non-regular Lagrangian concordances
@@
 Namely, these constructions produce exact Lagrangian cobordisms with an empty positive end inside a symplectisation; elementary topological considerations imply that such a cobordism cannot be regular.
-From this perspective, the example that we construct is different, it has a non-empty positive end and the simplest possible topology among Lagrangian cobordisms, namely that of a concordance.
-First we need the following generalisation of [7, Theorem 3.2]
+From this perspective, the examples that we construct here are different, since they have a non-empty positive end and the simplest possible topology among Lagrangian cobordisms; they are concordances.
+First we need the following generalisation of [7, Theorem 3.2].
 Theorem 5.1.
-Let $C\subset[-T,T]\times S^{3}$ be a smooth concordance that coincides with the cylinder over a knot $K_{\pm}$ near the boundary, and such that $[-T,T]\times S^{3}$ can be obtained by attaching $0,1,$ and $2$-handles onto a neighborhood of
-($(\{-T\}\times S^{3})\cup C\subset[-T,T]\times S^{3}.$
+Let $C\subset[-T,T]\times S^{3}$ be a smooth concordance that coincides with the cylinder over a knot $K_{\pm}$ near the boundary, and such that $T,T]\times S^{3}$ can be obtained by attaching $0,1,$ and $2$-handles onto a neighborhood of
+$(\{-T\}\times S^{3})\cup C\subset[-T,T]\times S^{3}.$
 If $K_{+}$ is the unknot, then so is $K_{-}$.
@@
 The proof is an adaptation of the proof of [7, Theorem 3.2], which showed that such a concordance $C$ cannot be ribbon, i.e. the symplectisation coordinate must have a local maximum somewhere.
-Consider the handle-decomposition of $[-T,T]\times S^{3}$ in which all handles are attached onto a neighbourhood ($(\{-T\}\times S^{3})\cup C$, and for which there are no $3$- or $4$-handles.
-We may also assume that there are no 00-handles, and that the $2$-handles are attached after the $1$-handles.
+Consider the handle-decomposition of $T,T]\times S^{3}$ in which all handles are attached onto a neighbourhood $(\{-T\}\times S^{3})\cup C$, and for which there are no $3$- or $4$-handles.
+We may also assume that there are no 0-handles, and that the $2$-handles are attached after the $1$-handles.
 Since $H_{*}([-T,T]\times S^{3},\{-T\}\times S^{3})=0$, it follows that the number of $1$-handles is equal to the number of $2$-handles; denote this number by $h$.
-[…]
+Then we consider the double cover of $T,T]\times S^{3}$ branched over the concordance $C$.
+This branched cover thus has a handle decomposition for which there are $2h$ handles in the cover of index one as well as of index two.
+This branched double cover is a cobordism from $\Sigma(K_{-})$ to $\Sigma(K_{+})$, where $\Sigma(K_{\pm})$ denotes the double cover of $S^{3}$ branched along the knot $K_{\pm}$.
+Note that $\Sigma(K_{+})=S^{3}$ since $K_{+}$ is the unknot.
 Attaching a four-handle to the end $\Sigma(K_{+})=S^{3}$ we obtain a cobordism $X$ from $\Sigma(K_{-})$ to $\emptyset$ which is simply connected.
-Indeed, turning it upside down, we have a cobordism with only 00-, $2$-, and $3$-handles.
-[…]
+Indeed, turning it upside down, we have a cobordism with only 0-, 22-, and $3$-handles.
+If we then consider $X$ as obtained by attaching $1$, 22, and $4$-handles on $\Sigma(K_{-})$, this means that the trivial group can obtained from $\pi_{1}(\Sigma(K_{-}))$ by adding $2h$ new generators (coming from the one-handles) and $2h$ new relations (coming from the two-handles).
 The rest of the proof is the same as the argument in [7, Theorem 3.2].
-In particular, this argument uses the adaptation [7, Theorem 3.1] of a result of Kronheimer and Mrowka that shows that that $\pi_{1}(\Sigma(K_{-}))$ admits a non-trivial representation $\rho$ in $SO(3)$ if and only if $K_{-}$ is not the unknot, while the argument in [29, Section I.3] shows that there can be no such representation under the assumption that $2h$ new generators and relations kill the group $\pi_{1}(\Sigma(K_{-}))$.
-[…]
+In particular, this argument uses the adaptation [7, Theorem 3.1] of a result of Kronheimer and Mrowka that shows that that $\pi_{1}(\Sigma(K_{-}))$ admits a non-trivial representation $\rho$ in $SO(3)$ if and only if $K_{-}$ is not the unknot, while the argument in [31, Section I.3] shows that there can be no such representation under the assumption that $2h$ new generators and relations kill the group $\pi_{1}(\Sigma(K_{-}))$.
+We finish by recalling the latter argument.
+The group $G_{0}\coloneqq\pi_{1}(\Sigma(K_{-}))/\ker\rho$ is a non-trivial group by the assumption that $\rho$ is non-trivial.
+This quotient group can be taken to have the same generators as $\pi_{1}(\Sigma(K_{-}))$ (in particular it is finitely generated).
+In addition, it admits a faithful non-trivial representation into $SO(3)$.
+Adjoining $2h$ new generators, i.e.
+considering the free product $G_{0}\star\mathbb{F}_{2h}$ of $G_{0}$ and the free group on $2h$ generators, together with the $2h$ relations as above, we obtain a group $G$.
+This group is the quotient of the free product $\pi_{1}(\Sigma(K_{-}))\star\mathbb{F}_{2h}$ by the normal subgroup generated by the $2h$ relations together with the relations from $\ker\rho$.
+In particular, $G$ is the quotient of the trivial group $\pi_{1}(X)$ and, thus, $G$ is itself trivial.
+Finally we can use [26, Theorem 3] to show that $G_{0}$ embeds in $G$, which means that $G_{0}$ is trivial, and thus leading to the sought contradiction.
+(Note that the condition on the determinant of the $2h$ relations needed in the cited theorem is satisfied here by the argument above, by which $H_{1}(\Sigma(K_{-}))$ is killed inside $H_{1}(X)$ by these $2h$ relations in the abelianisation.)
 Remark 5.2.
 It is unknown to the authors whether a regular Lagrangian cobordism is also decomposable.
-A decomposable Lagrangian cobordism $L\subset[-T,T]\times S^{3}$ is regular by a particular Weinstein handle decomposition of $[-T,T]\times S^{3}$ in which the handle decomposition of the neighbourhood $T^{*}L$ and the handles attached to $\{-T\}\times S^{3}\cup T^{*}L$ are in cancelling position.
-It is not clear if a general set of Weinstein handle attachment on $\{-T\}\times S^{3}\cup T^{*}L$ that produces $[-T,T]\times S^{3}$ can be deformed into such a position.
-In recent work [8] by the first author it is shown that totally real concordances can be made Lagrangian after adding sufficiently many stabilisations to the Legendrian ends.
+A decomposable Lagrangian cobordism $L\subset[-T,T]\times S^{3}$ is regular by a Weinstein handle decomposition of $T,T]\times S^{3}$ of a very particular type, in which the handle decomposition of the neighbourhood $T^{*}L$ and the handles attached to $\{-T\}\times S^{3}\cup T^{*}L$ are in cancelling position.
+It is not clear if a general set of Weinstein handle attachment on $\{-T\}\times S^{3}\cup T^{*}L$ that produces $T,T]\times S^{3}$ can be deformed into such a position.
+In recent work [8] by the first author it is shown that totally real concordances can be deformed in order to become Lagrangian after adding sufficiently many stabilisations to the Legendrian ends.
+To formulate the result we need to recall the following definition.
+An almost complex structure on the symplectisation $(\mathbb{R}_{t}\times Y,d(e^{t}\,\alpha))$ of a contact manifold $(Y,\alpha)$ is called cylindrical if
+  • $J$ is invariant under translation of the $t$-coordinate;
+  • $J$ preserves the contact planes $\ker\alpha\subset TY$ in each level-set $\{t\}\times Y$ and is compatible with $d\alpha$ inside the same, i.e.
+$d\alpha(\cdot,J\cdot)$ defines a Riemannian metric on $\ker\alpha$; and
+  • $J\partial_{t}=R_{\alpha}$ where $R_{\alpha}$ is the Reeb vector field.
 Theorem 5.3 ([8]).
-Assume that there exists a totally real concordance from a Legendrian knot $\Lambda_{-}^{\prime}$ to $\Lambda_{+}^{\prime}$ (which according to Appendix A implies that $\Lambda_{\pm}^{\prime}$ are smoothly concordant, and with the same ${\tt tb}$ and ${\tt rot}$).
+Assume that for a given choice of cylindrical almost complex structure on $(\mathbb{R}\times S^{3},{\color[rgb]{0,0,0}d(e^{t}\,\alpha_{st})})$ there exists a totally real concordance from a Legendrian knot $\Lambda_{-}^{\prime}$ to $\Lambda_{+}^{\prime}$ (which according to Appendix A implies that $\Lambda_{\pm}^{\prime}$ are smoothly concordant, and with the same $\tt tb$} and $\tt rot$}).
 Then this concordance can be deformed to a Lagrangian concordance from $\Lambda_{-}$ to $\Lambda_{+}$ by a smooth isotopy, where both $\Lambda_{\pm}$ are obtained from $\Lambda_{\pm}^{\prime}$ by performing a sufficiently large number $k\gg 0$ of stabilisations of both signs.
@@
 Proposition 5.4.
-If $C\subset\mathbb{R}_{t}\times\mathbb{R}^{3}$ is a decomposable Lagrangian concordance from $\Lambda_{-}$ to $\Lambda_{+}$ with $\tt{rot}(\Lambda_{\pm})=0$, then the reversed cobordism
+For any decomposable Lagrangian concordance
+$C\subset(\mathbb{R}_{t}\times\color[rgb]{0,0,0}\mathbb{R}^{3}_{xyz},d(e^{t}(dz-y\,dx)))$
+from a Legendrian knot $\Lambda_{-}$ to a Legendrian knot $\Lambda_{+}$, the reversed cobordism
 $C^{\prime}\coloneqq\{(t,x,y,z)\in\mathbb{R}_{t}\times\mathbb{R}^{3};\>(-t,x,y,z)\in C\}$
-admits a compactly supported smooth isotopy to a concordance that is totally real for any choice of cylindrical almost complex structure.
+admits a compactly supported smooth isotopy to a cobordism that is totally real, for any given choice of cylindrical almost complex structure.
 Proof.
-First, recall that the space of oriented real two-planes $V\subset\mathbb{R}^{4}$ can be identified with $\mathbb{C}P^{1}\times\mathbb{C}P^{1}\cong Gr^{+}_{2}(\mathbb{R}^{4})$.
-Here we identify the first $\mathbb{C}P^{1}$-factor with an element in the twistor-space $SO(4)/U(2)$, i.e.
-an almost complex structure $J\in SO(4)$, while the second factor denotes a $J$-complex line in $\mathbb{R}^{4}$.
-For an appropriate such identification, the Lagrangian planes are given by the subset
-$S^{1}\times\mathbb{C}P^{1}\subset\mathbb{C}P^{1}\times\mathbb{C}P^{1}\cong Gr^{+}_{2}(\mathbb{R}^{4}),$
-while the non-totally real 2-planes for the standard complex structure $J_{0}$ can be identified with
-$D=\{0,\infty\}\times\mathbb{C}P^{1}\subset\mathbb{C}P^{1}\times\mathbb{C}P^{1}\cong Gr^{+}_{2}(\mathbb{R}^{4}).$
-More precisely, the latter submanifold consists of those planes that coincide with $J_{0}$-complex planes equipped either with the complex or anti-complex orientation.
-It follows from the above that any smooth concordance $C^{\prime}$ between $\Lambda_{\pm}$, where $\tt{rot}(\Lambda_{\pm})=0$, that is cylindrical outside of a compact subset has a Gauss map of tangent planes
-$\Gamma\colon C^{\prime}\to Gr^{+}_{2}(\mathbb{R}^{4})$
-that is compactly supported homotopic to a family of totally real planes if and only if the intersection number with the hypersurface $D\subset\mathbb{C}P^{1}\times\mathbb{C}P^{1}\cong Gr^{+}_{2}(\mathbb{R}^{4})$ vanishes.
-Indeed, the assumptions on the boundary imply that the Gauss map along $\partial C^{\prime}$ maps to a null-homologous curve inside the equator $S^{1}\subset\mathbb{C}P^{1}=SO(4)/SU(2)$ when projected to the first factor in the above identification $\mathbb{C}P^{1}\times\mathbb{C}P^{1}\cong Gr^{+}_{2}(\mathbb{R}^{4})$.
-Thus, the algebraic number of pre-images of the two different components
-$\{0\}\times\mathbb{C}P^{1}\>\sqcup\>\{\infty\}\times\mathbb{C}P^{1}=D$
-under the Gauss map can readily be seen to agree.
-When both of these two numbers vanish (which, under the current assumptions, is the same as the vanishing of the total algebraic number of pre-images of $D$), one can readily find a homotopy of the Gauss map relative boundary that removes all intersections with $D$.
-By the $h$-principle for totally real embeddings relative to boundary [20, 27.4.1 and Example 2], we simply need to show that $C^{\prime}$ has a Gauss map that is homotopy relative boundary into the totally real planes in order to ensure the existence of the sought smooth isotopy.
-Using the argument in the previous paragraph, it thus only remains to show that the reversed cobordism $C^{\prime}$ has a Gauss map for which the count of pre-images of $D$ vanishes algebraically.
-Since $C$ is decomposable, this follows from a local computation in the standard model.
-Indeed, by the definition of decomposability, $C$ can be written as a concatenation of Lagrangian traces of Legendrian isotopies and Lagrangian standard 00 and $1$-handles.
-The subset of $C^{\prime}$ that corresponds to a Lagrangian trace is clearly smoothly isotopic to a totally real cobordism, since it is even smoothly isotopic to the Lagrangian trace of the inverse isotopy.
-It thus suffices to compute the intersections with $D$ for a standard Lagrangian 00 and $1$-handle.
-Here, note that away from a small neighbourhood of the critical point of the standard Lagrangian handle, we can assume that the Lagrangian is cylindrical and thus the reversed cobordism is again Lagrangian.
-Since $C$ is a cylinder there is an equal number of 00 and $1$-handles.
-It can be seen that these two types of handles contribute with intersection numbers with $D$ of opposite signs; indeed, this can be checked by putting two Lagrangian 00 and $1$-handles in cancelling position.
+Outside of a sufficiently large compact subset that contains the non-cylindrical part of $C$, the reversed cobordism $C^{\prime}$ is already Lagrangian, and thus totally real for any compatible almost complex structure.
+We will analyse the homotopy class of the tangent planes of $C^{\prime}$ inside $\mathbb{R}\times\mathbb{R}^{3}$, in order to show that the map is compactly supported homotopic to one which takes values in Lagrangian tangent planes; the latter are in particular totally real for any compatible almost complex structure.
+The existence of the smooth isotopy is then a consequence of the $h$-principle for totally real embeddings relative to the boundary [21, 27.4.1 and Example 2].
+The assumption of decomposability implies that the concordance consists of $k$ number of standard 0-handles and $k$ number of standard 1-handles of the types described in Section 3.1.
+We may assume that all 0-handles are attached before the 1-handles.
+The 0-handles give rise to a number $k$ of small unlinked unknots.
+Instead of applying Legendrian isotopies to these unknots and $\Lambda_{-}$, and then performing the cusp-connected sums, we are able to fix both $\Lambda_{-}$ and the $k$ unknots, and then attach the 1-handles along long and complicated (e.g.
+possibly non-trivially linked) attaching arcs; see Subsections 3.1 and 3.3.
+An example of this procedure is shown in Figure 18.
+The decomposable concordance can thus be described by starting with the Legendrian $\Lambda_{-}$ and $k$ unlinked unknots, together with $k$ complicated surgery arcs.
+Since the cobordism is a connected concordance, it follows that the surgery arcs corresponding to the 1-handles must connect the $k$ unlinked unknots and the negative end $\Lambda_{-}$; the graph whose $k+1$ vertices correspond to the Legendrian knots and $k$ edges corresponding to the surgery arcs is thus a connected tree.
+After performing a Legendrian regular homotopy (not Legendrian isotopy!) of the surgery arcs, we may assume that the surgery arcs all become unlinked.
+To see this it suffices to apply the $h$-principle for Legendrian immersions, but one can also see this by an explicit deformation using the front projection.
+An example of such a regular homotopy is the one that takes the non-trivially linked surgery arc shown in Figure 18 and replaces it with the unlinked surgery arc shown in Figure 19.
+Consequently, this means that $C$ admits a regular homotopy through Lagrangians that are
+  • coinciding with $(-\infty,-A]\times\Lambda_{-}$ in $(-\infty,-A]\times\mathbb{R}^{3}$; and
+  • cylindrical (and possibly immersed) in $A,+\infty)\times\mathbb{R}^{3}$,
+for $A\gg 0$, to an embedded decomposable Lagrangian concordance $\tilde{C}$ for which the $k$ number of Lagrangian 0-handles can be cancelled with the $k$ number of Lagrangian 1-handles.
+In other words $\tilde{C}$ is Hamiltonian isotopic to a trivial concordance.
+We will only need the following weaker consequence of the above regular homotopy: the Lagrangian gauss map of $C$ admits a compactly supported homotopy to a map that takes values in Lagrangian planes that are everywhere tangent to $\partial_{t}$, i.e.
+that takes values in cylindrical Lagrangian tangent planes.
+Note that the smooth involution $(t,(x,y,z))\mapsto(-t,(x,y,z))$ of the symplectisation $\mathbb{R}\times\mathbb{R}^{3}$ preserves those two-planes that are tangent to $\partial_{t}$, e.g.
+the cylindrical Lagrangian tangent planes.
+Since $C$ has a Gauss map that is compactly supported homotopic to a map of cylindrical Lagrangian planes, we conclude that the tangent map of $C^{\prime}$ is also compactly supported homotopic to a map of Lagrangian tangent planes.
+This shows that the $h$-principle for totally real embeddings can be applied to $C^{\prime}$ as sought.
 ∎
 Remark 5.5.
-Note that totally real concordances define a relation which is obviously reflexive and transitive.
+Note that totally real concordances define a relation on the set of Legendrian isotopy classes of Legendrian knots which is obviously reflexive and transitive.
 From Proposition 5.4 it follows that this relation is not anti-symmetric, and hence totally real concordances do not define a partial order.
@@
 By Theorem 5.1 the Lagrangian concordances produced by Corollary 5.6 are never regular.
-Therefore, from the result of Conway, Etnyre and Tosun [6], it follows that these Lagrangian concordances also are not decomposable in the sense of Chantraine [4].
+Therefore, from the result of Conway, Etnyre and Tosun [6] (or, alternatively, from our generalisation Proposition 3.2), it follows that these Lagrangian concordances also are not decomposable in the sense of Chantraine [4].
 Proof.
-A decomposable Lagrangian disc filling can be seen as a standard disc filling of a standard Legendrian unknot $U$ concatenated with a decomposable concordance from $U$ to $\Lambda$.
-We apply Proposition 5.4 produce a totally real concordance from $\Lambda$ to $U$.
-Theorem 5.3 then produces the sought Lagrangian concordance, by which we first need to add sufficiently many positive and negative stabilisations of to $U$ and $\Lambda$.
+A decomposable Lagrangian disc filling can be seen as a standard disc filling of a standard Legendrian unknot $U$ concatenated with a concordance from $U$ to $\Lambda$.
+We apply Proposition 5.4 in order to produce a totally real concordance from $\Lambda$ to $U$.
+Theorem 5.3 then gives us the sought Lagrangian concordance, by which we first need to add sufficiently many positive and negative stabilisations to $U$ and to $\Lambda$.
 ∎
@@
 Theorem A.1.
-If there is a totally real orientable cobordism $L$ from Legendrian knot $\Lambda_{-}$ to Legendrian knot $\Lambda_{+}$, then
-$tb(\Lambda_{+})-tb(\Lambda_{-})=-\chi(\Lambda),$
+If there is a totally real orientable cobordism $L$ from null-homologous Legendrian knot $\Lambda_{-}$ with a Seifert surface $S$ to Legendrian knot $\Lambda_{+}$, then
+$tb(\Lambda_{+})-tb(\Lambda_{-})={\color[rgb]{0,0,0}-\chi(L)},$
 $r(\Lambda_{-},[S])=r(\Lambda_{+},[S\cup L]).$
@@
 The proof is completely analogous to the case when the cobordism is Lagrangian.
-Namely, we perturb the symplectisation coordinate in a compact subset so that it restricts to a Morse function $h\colon S\to\mathbb{R}$ on the cobordism.
+Namely, we perturb the symplectisation coordinate in a compact subset so that it restricts to a Morse function $h\colon\color[rgb]{0,0,0}L\color[rgb]{0,0,0}\to\mathbb{R}$ on the cobordism.
 Then we consider a push-off by $J(\nabla h)$ for an almost complex structure that is cylindrical outside of a compact subset, compatible with the symplectic structure, and for which the cobordism is totally-real.
 We can assume that $\nabla h=\partial_{t}$ is satisfied outside of a compact subset.
-It follows that the signed count of intersection points are precisely $-\chi(S)$ (as in the Lagrangian case).
+It follows that the signed count of intersection points are precisely $\chi(\color[rgb]{0,0,0}L)$ (as in the Lagrangian case).
 The rest follows in the same manner.